### v1

# Generative Modeling with Explicit Memory

Abstract
Corresponding author.* Equal contribution. Peng is the project leader.
Recent studies indicate that the denoising process in deep generative diffusion models implicitly learns and memorizes semantic information from the data distribution. These findings suggest that capturing more complex data distributions requires larger neural networks, leading to a substantial increase in computational demands, which in turn become the primary bottleneck in both training and inference of diffusion models. To this end, we introduce Generative Modeling with Explicit Memory (GMem), leveraging an external memory bank in both training and sampling phases of diffusion models. This approach preserves semantic information from data distributions, reducing reliance on neural network capacity for learning and generalizing across diverse datasets. The results are significant: our GMem enhances both training, sampling efficiency, and generation quality. For instance, on ImageNet at $256\times 256$ resolution, GMem accelerates SiT training by over 46.7×46.7$\times$46.7 ×, achieving the performance of a SiT model trained for $7M$ steps in fewer than $150K$ steps. Compared to the most efficient existing method, REPA [70], GMem still offers a 16×16$\times$16 × speedup, attaining an FID score of 5.75 within $250K$ steps, whereas REPA requires over $4M$ steps. Additionally, our method achieves state-of-the-art generation quality, with an FID score of 3.56 without classifier-free guidance on ImageNet $256\times 256$. Our code is available at https://github.com/LINs-lab/GMem.
Figure 1: GMem significantly enhances the training and sampling efficiency of diffusion models on ImageNet $256\times 256$. We propose an external memory bank that enhances the semantic information storage, improving both training stability and sampling efficiency in diffusion models. Sub-figure (a) illustrates the key concept of GMem, where $\boldsymbol{\epsilon}$ represents the initial noise and $\mathbf{x}_{0}$ the generated samples. In GMem, we disentangle the generalization and memorization capabilities originally intertwined within diffusion models, assigning memorization to an external memory bank $\mathbf{M}$. This decoupling reduces computational and capacity overhead for memorizing internal semantic information, thereby accelerating the overall process. Sub-figure (b) demonstrates the training efficiency of GMem on ImageNet $256\times 256$. At an FID of $5.8$, GMem achieves over 16×16$\times$16 × speedup relative to REPA [70]. At an FID of $7.3$, GMem achieves over 46.7×46.7$\times$46.7 × speedup relative to SiT [44]. Sub-figure (c) illustrates the sampling efficiency of GMem on ImageNet $256\times 256$. At the same target level of FID, GMem requires 5×5$\times$5 × less NFEs compared to REPA and 10×10$\times$10 × less NFEs compared to SiT.
1 Introduction
Deep generative models based on denoising [68, 24, 59, 47, 9, 40, 8, 37]—which prioritize the generation of high-quality, realistic data—have achieved notable success within the deep learning community. These methods demonstrate exceptional performance in complex tasks such as zero-shot text-to-image and video generation [50, 56, 15, 51, 4]. However, training and sampling diffusion models suffer from the issue of high computational burden [28, 29], due to the necessity of using neural networks with larger capacities for better empirical performance [28, 29].
In the meanwhile, a line of research examines the diffusion models from the perspective of representation learning [36, 66, 8, 46], where diffusion models can capture semantic features within their hidden layers, with advanced models producing even stronger representations [66]. Yang and Wang [68] illustrate that diffusion models strike a balance between learning meaningful features and regularizing the model capacity. Kadkhodaie et al. [27] further interpret the generalization behavior in diffusion models, indicating that diffusion models aim to learn and encapsulate semantic information from the training data, subsequently generalizing to the true data distribution.
In this paper, we extend these insights [68, 27] and conjecture that the functions of diffusion models can be implicitly divided into two components: (i) memorization of semantic information and (ii) generalization to true data. The intrinsic generation procedure of diffusion model then can be explained as transforming input noisy samples into the internal semantic distribution of a model, followed by “decoding” it into the data distribution. Specifically,
  1. (a)
    in the training stage, diffusion models require significant neural network parameters and computational resources to extract and memorize semantic information from noisy samples;
  2. (b)
    in the sampling/inference stage, models often remap Gaussian noise to the internal semantic distribution before retrieving the data distribution, rather than directly transforming noise to data.
Building upon the aforementioned analyses, the efficiency challenges of training and sampling diffusion models stem from the need to learn and memorize the internal semantic information, resulting in higher computational cost and greater demands on model capacity. As a remedy, we propose GMem, a new paradigm that uses an external memory bank to accompany the diffusion models to approximate the true data distribution, therefore avoiding memorizing the data distributions in the neural network. Extensive experiments demonstrate the effectiveness and efficiency of our approach, on both diffusion transformers [44] and convolutional neural networks [24, 61, 28]. For example, integrating our GMem with SiT on ImageNet $256\times 256$ only requires $150K$ training iterations to achieve an FID of $7.3$, which is 46.7×46.7$\times$46.7 × faster compared to the baseline SiTs [44] and 16×16$\times$16 × than REPA [70]. Under the same number of training iterations as $400K$, our model enhances the sampling efficiency by 10×10$\times$10 ×, where our model requires only $25$ sampling steps to achieve an $\texttt{FID}=12.3$ while the vanilla SiT needs $250$ sampling steps. Ultimately, our approach achieves SoTA performance with an FID of $3.56$ after $2M$ training steps, without classifier-free guidance.
We outline the main contributions of this paper as follows:
  1. (a)
    We propose that diffusion models operate through two fundamental perspectives: memorization and generalization. Utilizing neural networks for memorization incurs significant computational and model capacity burdens.
  2. (b)
    We present GMem, a simple yet extremely effective approach to construct an external memory bank that stores semantic information inherent in data distributions, alleviating the memorization burden on neural networks within the diffusion models.
  3. (c)
    Our framework significantly enhances efficiency and quality of the SoTA diffusion models, for both training and inference (see Figure 1 and Figure 2).
Figure 2: Selected samples on ImageNet $256\times 256$. This figure showcases a selection of images generated by GMem. We utilized the SiT-XL/2 backbone and trained GMem for $2M$ steps without classifier-free guidance.
2 Related work
Generative models.
Generative models—including Generative Adversarial Networks (GANs) [19, 57, 67], Variational Autoencoders (VAEs) [32, 21], flow-based methods, and diffusion-based methods [24, 12, 45]—aim to learn the data distribution $p(\mathbf{x})$ and generate data through sampling, demonstrating remarkable performance in producing realistic samples [38]. Recently, diffusion-based methods utilize stochastic interpolation to model the forward process and subsequently convert the Gaussian distribution back into the original image distribution during the reverse process, yielding realistic samples. Such methods have achieved SoTA results in deep generative modeling and are the focus of this study [45, 60, 14].
Training diffusion-based models is computationally intensive and unstable [70, 60]. Additionally, the generative process necessitates multi-step sampling, incurring substantial computational costs [42] Accelerating the training and sampling processes of diffusion models has emerged as a significant challenge, prompting various efforts in previous studies to address this issue.
Diffusion modeling and representation learning.
To overcome the instability and computational inefficiency of diffusion models, recent studies  [70, 17, 45] have turned to representation learning to enhance diffusion models. On one hand, diffusion models are capable of learning high-quality representations [70]. For instance, Tang et al. [63] demonstrated that feature maps extracted from diffusion networks can establish correspondences between real images, indicating a strong alignment between the learned representations with actual image. Furthermore, Yang and Wang [68] conducted a detailed analysis of the trade-off between the quality of learned representations and the penalization of the optimal parameter spectrum.
Conversely, well-trained representation models can improve performance and expedite the training of diffusion models. [45] accomplish this by adjusting the weighting function in the denoising score matching objective to enhance representation learning. REPA [70] introduces an alignment loss for intermediate layer representations, significantly accelerating the training process by over $17.5$ times. In general, these methods enhance diffusion modeling by employing regularization techniques, which implicitly integrate semantic knowledge from pre-trained representation models.
Self-supervised learning in diffusion models.
Self-supervised learning, as a powerful approach within representation learning that leverages unlabeled data to learn useful representations, has garnered increasing attention in the diffusion modeling community. Many recent works [39, 26, 2] have adopted self-supervised learning techniques to enhance the representation learning capabilities of diffusion models, thereby improving their performance and efficiency. In particular, Li et al. [39] proposed a novel diffusion training paradigm that introduces an additional diffusion model to generate a 1D feature map as a condition to guide the training of the generative model. However, the inclusion of an additional diffusion model introduces extra sampling during inference, which increases computational overhead and can introduce additional errors. In this work, differing from Li et al. [39], we directly utilize representation from ‘memory bank’ as condition. This straightforward strategy eliminates the need for an additional diffusion model, thereby enhancing representation learning while reducing computational costs. It significantly improves both the training speed and the effectiveness of the model. In this work, differing from Li et al. [39], we directly utilize representation from ‘memory bank’ as condition. This straightforward strategy not only eliminates the unnecessary computational burden but also serves as an effective self-supervised learning method that enabling the model to learn better representations. As a result, our method significantly enhances both the training speed and effectiveness of diffusion model.
Text to image generation.
Text-to-image generation seeks to synthesize realistic images from free-form textual prompts. Despite advancements, current text-to-image models continue to experience challenges with text-image misalignment [53, 35]. Recent works [6, 16] have attempted to address this issue by intervening in cross-modal attention activations that influence the internal states of Stable Diffusion models. However, this approach increases computational complexity and may impede the learning of optimal representations. Indeed, these models often require several days to weeks for training.
3 Preliminary
In this section, we provide a concise introduction to the training and sampling processes of flow-based and diffusion-based models. See Appendix G for more details.
Both diffusion-based and flow-based models [44] derive their training procedures from a deterministic $T$-step noising process applied to the original data, formalized as:
$\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})\,,$ (1)
where $\mathbf{x}_{t}$ represents the noisy data at time $t$, $\mathbf{x}_{0}\sim p(\mathbf{x})$ is a real data sample from the true distribution, $\alpha_{t}$ and $\sigma_{t}$ are time-dependent decreasing and increasing functions respectively satisfying $\alpha_{t}^{2}+\sigma_{t}^{2}=1$.
According to (1), each marginal probability density $p_{t}(\mathbf{x}_{t})$ corresponds to the distribution of a Probability Flow Ordinary Differential Equation [61] (PF ODE) with velocity field $\mathbf{v}(\mathbf{x},t)$, defined as:
$\mathbf{v}(\mathbf{x},t)=\dot{\alpha}_{t}\,\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t}=\mathbf{x}]-\dot{\sigma}_{t}\,\mathbb{E}[\boldsymbol{\epsilon}\mid\mathbf{x}_{t}=\mathbf{x}]\,,$ (2)
where $\dot{\alpha}_{t}=\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}$ and $\dot{\sigma}_{t}=\frac{\mathrm{d}\sigma_{t}}{\mathrm{d}t}$. By solving this ODE starting from $\mathbf{x}_{T}=\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, we obtain the probability density function $p_{0}(\mathbf{x}_{0})$, which can be used to estimate the ground-truth data distribution $p(\mathbf{x})$.
Alternatively, the aforementioned noise-adding process can be formalized as a Stochastic Differential Equation [61, 59] (SDE):
$\mathrm{d}\mathbf{x}_{t}=\boldsymbol{m}(\mathbf{x}_{t},t)\,\mathrm{d}t+g(t)\,\mathrm{d}\mathbf{W}_{t}\,,$ (3)
where $\mathbf{W}_{t}$ is a Wiener process [23], $\boldsymbol{m}(\mathbf{x}_{t},t)$ is the drift coefficient defined as $\boldsymbol{m}(\mathbf{x}_{t},t)=-\frac{1}{2}\beta(t)\mathbf{x}_{t}$, and $g(t)$ is the diffusion coefficient, set as $g(t)=\sqrt{\beta(t)}$ with $\beta(t)$ being a time-dependent positive function controlling the noise schedule.
The corresponding reverse process is represented by the reverse-time SDE:
$\mathrm{d}\mathbf{x}_{t}=\left[\boldsymbol{m}(\mathbf{x}_{t},t)-g(t)^{2}s(\mathbf{x}_{t},t)\right]\mathrm{d}t+g(t)\,\mathrm{d}\bar{\mathbf{W}}_{t}\,,$ (4)
where $\bar{\mathbf{W}}_{t}$ is a reverse-time Wiener process, and $s(\mathbf{x}_{t},t)$ is the score function, defined by the gradient of the log probability density:
$s(\mathbf{x}_{t},t)=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})=-\frac{1}{\sigma_{t}}\mathbb{E}\left[\boldsymbol{\epsilon}\mid\mathbf{x}_{t}=\mathbf{x}\right]\,.$ (5)
By solving the reverse-time SDE (4), starting from the initial state $\mathbf{x}_{T}=\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, we can obtain $p_{0}(\mathbf{x}_{0})$, thereby estimating the true data distribution $p(\mathbf{x})$.
4 Methodology
Algorithm 1 Training GMem using memory bank $\mathbf{M}$
procedure Train GMem($\mathbf{v}_{\boldsymbol{\theta}},\mathcal{D},\mathbf{M},T,\alpha_{t},\sigma_{t}$)
     Initialize model parameters $\boldsymbol{\theta}$
     for each training iteration do
         Sample a batch of data $\mathbf{x}_{0}\sim\mathcal{D}$
         Sample timesteps $t\sim\{0,\dots,T\}$ uniformly
         Generate noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$
         Compute noisy data $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\boldsymbol{\epsilon}$
         Sample memory snippet $\mathbf{s}\sim\mathbf{M}$ using (9)
         Predict velocity $\mathbf{v}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\mathbf{s})$
         Compute loss using (8) and (10)
         Backpropagate and update $\boldsymbol{\theta}$ using Optimizer
     end for
end procedure
The motivation for our method arises from the observation that diffusion models inherently encode two distinct capabilities: generalization and memorization of semantic information within data [70, 68, 66, 8]. We propose a paradigm in diffusion models that separates generalization and memorization: neural networks are employed for generalization, while an external memory bank is used to store semantic information of data.
Specifically, we find that conventional generative models [24, 60] typically use a neural network to achieve both generalization and memorization of true data distributions. However, the UNet and Transformer-based neural networks [24, 49] used in modern diffusion models have limited parameters. This makes it challenging for these networks to memorize distribution information, leading to two main issues [27]: (a) limited parameters reduce the ability to memorize extensive data distributions, and (b) memorization via parameter optimization incurs high computational and time costs.
To address these challenges and alleviate the memorization burden on neural networks in diffusion models, we propose employing an external memory bank to serve as the memorization function, therefore significantly improving training efficiency and the models’ ability to capture distributions. The framework is depicted in Figure 3.
4.1 External Memory Bank
Inspired by previous studies [27], which demonstrated that generalization in diffusion models arises from geometry-adaptive harmonic representations, we aim to design a memory bank that: (a) Provides sufficient semantic and representational information to assist the diffusion model in generating high-quality, realistic images that align with the true data distribution. (b) Avoids excessive detailed or textual information to prevent overfitting to the training data and ensure robust generalization.
We define the memory bank for diffusion models as a matrix $\mathbf{M}\in\mathbb{R}^{n\times m}$, encompassing $n$ memory snippets. Each snippet $\mathbf{s}$ contains $m$ units and satisfies $\|\mathbf{s}\|=1$:
$\mathbf{M}=\begin{bmatrix}\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{n}\end{bmatrix}^{\top}\,.$ (6)
We employ a robust representation model $\boldsymbol{f}$ under the condition that for every input $\mathbf{x}$ sampled from data distribution $D$, the produced normalized feature $\nicefrac{{\boldsymbol{f}(\mathbf{x})}}{{\|\boldsymbol{f}(\mathbf{x})\|}}$ always corresponds to a vector $\mathbf{s}$ that exists in the set $\mathbf{M}$. In conclusion, our goal is to ensure that all semantic information contained in $D$ is captured by $\mathbf{M}$.
To achieve this goal, we formalize the optimization objective for $\mathbf{M}$ as follows (see our analysis in Appendix A.1):
$\operatorname*{arg\,min\,}_{\mathbf{M}\in\mathbb{R}^{n\times m}}\left\{\left(\mathbb{E}_{\mathbf{x}\sim D}\left[\max\left\{\mathbf{M}\frac{\boldsymbol{f}(\mathbf{x})}{\|\boldsymbol{f}(\mathbf{x})\|}\right\}-1\right]\right)^{2}\right\}\,,$ (7)
where $\max(\cdot)$ denotes the maximum value of a vector. Note that $n$ and $m$ are predefined hyperparameters balancing the memorization capacity of $\mathbf{M}$ against storage constraints. Specifically, increasing $n$ and $m$ enhances the memorization capacity of $\mathbf{M}$ but also increases storage requirements. Detailed settings and analyses of $n$ and $m$ are provided in Appendix A.2.
Usage of representation models for building the memory bank.
Our memory bank framework is adaptable to various representation models $\boldsymbol{f}$, each with distinct capabilities in capturing image information. This variability impacts the performance of the trained generative models. Therefore, in this study, we utilize self-supervised representation models based on the following reasoning:
  1. (a)
    Previous research [70] demonstrates that self-supervised models can expedite the training of diffusion models, which can be seamlessly integrated with a memory bank.
  2. (b)
    Prior work [3, 74, 62] points out that self-supervised models capture semantic information more effectively than supervised alternatives.
We rigorously assess the effectiveness of different representation models within our framework to build the memory bank like CLIP visual encoder [54] and BERT [11] (see Appendix A.3 and Appendix F).
Figure 3: Data generation via GMem-enhanced diffusion models. (a) Sampled noise $\boldsymbol{\epsilon}$ is used to index a memory snippet from the memory bank. (b) Both the sampled noise $\boldsymbol{\epsilon}$ and the memory snippet $\mathbf{s}$ are simultaneously fed into the neural network. (c) The neural network generates data using SDE or ODE solvers.
4.2 Training Objective
Our proposed memory bank is designed to supply semantic information about the data distribution, supporting both the training and inference phases of diffusion models. Therefore, we incorporate semantic information into diffusion models during the training phase by slightly modifying the diffusion model loss function as follows:
[…]
where $\mathbf{x}_{0}\sim D,\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),\dot{\alpha}_{t}=\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t},\dot{\sigma}_{t}=\frac{\mathrm{d}\sigma_{t}}{\mathrm{d}t},\lambda\sim p(\lambda)$ and $\mathbf{v}_{\boldsymbol{\theta}}$ is the velocity estimated by SiT backbone. This approach trains diffusion models using a interpolant way over various noise levels [44], incorporating an additional input snippet $\mathbf{s}$, which was selected from memory bank $\mathbf{M}$ using cosine similarity:
$\mathbf{s}=\operatorname*{arg\,max\,}_{\mathbf{s}_{i}\in\mathbf{M}}\left\{\frac{\boldsymbol{f}(\mathbf{x}_{0})}{\|\boldsymbol{f}(\mathbf{x}_{0})\|}\cdot\mathbf{s}_{i}\right\}\,.$ (9)
To ensure a fair comparison with the baselines, we use the exact same parameter settings of $\alpha_{t}$ and $\sigma_{t}$ as in previous studies [44, 70]. For a detailed description of the training and generation processes, please refer to the framework outlined in Algorithm 1.
Multiple usage of representation models for enhancing training diffusion models.
A very recent work REPA [70] introduces a regularization term that leverages representation models within the diffusion model’s loss function to enhance training efficiency. Therefore, we can integrate this regularization with our modified diffusion loss (8):
$\mathcal{L}(\boldsymbol{\theta})\leftarrow\mathcal{L}(\boldsymbol{\theta})+\mathcal{L}_{\textsc{REPA}}(\boldsymbol{\theta})$ (10)
$\text{with}\;\;\mathcal{L}_{\textsc{REPA}}(\boldsymbol{\theta}):=-\mathbb{E}_{\mathbf{x}_{0},t}$ $\left[\operatorname{sim}\left(\boldsymbol{f}(\mathbf{x}_{0}),\boldsymbol{\psi}(\mathbf{h})\right)\right]\,,$
where $\boldsymbol{f}(\mathbf{x}_{0})$ denotes the pretrained representation, and $\boldsymbol{\psi}(\mathbf{h})$ represents the projection of the intermediate output of diffusion transformer block $\mathbf{h}$ using a trainable projection head $\boldsymbol{\psi}$, implemented as a multilayer perceptron (MLP).
4.3 Sampling with Memory Bank
The sampling stage of unconditional generation in diffusion models [24, 42] aims to transform Gaussian noise into the true data distribution within an end-to-end framework. This iterative process employs SDE or ODE solvers (refer to Section 3) to progressively convert the noise into the desired data distribution. Our framework GMem with the memory bank also follows this end-to-end generation process, namely, feeding the noise into the neural network, together with the corresponding memory snippet $\mathbf{s}$.
Specifically, our GMem transforms the noise distribution into a corresponding bank snippet index distribution (a.k.a., noise-to-index transformation) and samples indices from this distribution to obtain memory snippets. We propose a two-step transformation methodology below—namely (1) normal distribution to uniform distribution and (2) uniform distribution to index distribution—that rigorously converts Gaussian-distributed noise into the index distribution. Such a strategy allows the index distribution exhibiting two key characteristics: firstly, similar noise inputs should correspond to similar indices, thereby maintaining coherence in the selection process; secondly, each generated index should possess an equal probability of being selected during the training phase, promoting unbiased learning.
Step 1: normal distribution to uniform distribution.
The initial step leverages the Probability Integral Transform [52] to convert Gaussian-distributed noise into a uniformly distributed variable. Specifically, we begin with a high-dimensional noise vector $\boldsymbol{\epsilon}\in\mathbb{R}^{d}$ that follows a Gaussian distribution. By computing and scaling the sum of the vector, we obtain a one-dimensional Gaussian variable $\overline{\epsilon}=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}\epsilon_{i}$, which follows the standard normal distribution $\mathcal{N}(0,1)$. The cumulative distribution function of this Gaussian variable, denoted by $\Phi(\overline{\epsilon})$, transforms the Gaussian noise into a uniform distribution over the interval $[0,1]$ and can be mathematically expressed as:
[…]
Step 2: uniform distribution to index distribution.
Given that the memory bank size may be much smaller than the total number of images in the dataset, a direct uniform distribution may not accurately reflect the true index distribution. To address this, we first record the index distribution from the training set, denoted as $\mathbf{P}_{\text{index}}=[p_{1},\ldots,p_{i},\ldots,p_{n}]$, where each $p_{i}$ represents the probability of selecting the $i$-th snippet $\mathbf{s}_{i}$ during training. Utilizing inverse transform sampling [52] , we map the uniform variable to the index distribution by defining the variable $k$ as the smallest index $i$ for which the cumulative probability $\sum_{j=1}^{i}p_{j}$ exceeds the uniform variable:
$k=\min\left\{i\in\{1,\ldots,n\}\ \bigg{|}\ \Phi(\overline{\epsilon})\leq\sum_{j=1}^{i}p_{j}\right\}\,.$
Upon completion of aforementioned noise-to-index transformation, the original noise vector $\boldsymbol{\epsilon}$ and the selected memory snippet $\mathbf{s}_{k}$ are combined and fed into the diffusion transformer backbone. The backbone is then interfaces with ODE or SDE solvers [24, 59] to generate images.
5 Experiments
5.1 Experimental Setting
We list the experimental settings below (see more details in Appendix B).
Datasets.
For pixel space generation, we evaluate GMem on CIFAR-10 [34] for its diversity in classes and widespread use in benchmarking image generation models. We also evaluate the scalability of GMem on ImageNet $64\times 64$ [10] to assess the performance on larger resolution generation. We also conduct experiments on ImageNet $256\times 256$ [10] to evaluate the ability of GMem to model the latent space distribution, a key approach in current image generation research [28, 70, 44, 49].
Backbones.
Similar to prior image generation works [70], we utilizes SiT [44] as backbone. This architecture represent most recent advances in diffusion and flow-based transformers [70], and are widely used in image generation tasks [20, 7, 70].
Baselines.
For a fair comparison, we compare to the SoTA image generation methods on both training efficiency and performance. Specifically, for pixel-space image generation, we consider the following three categories of baselines: First, we compare GMem with traditional generative models, including Diffusion GAN [67], Diffusion StyleGAN [65], DMD2 [69]. We also compare the SoTA diffusion models with UNets, including DDPM [24], Score SDE [61], EDM [28], DPM-Solver [43], ADM [12], EDMv2 [29], CTM [30], SiD [72], EMD [28]. Finally, we also compare to the SoTA flow-based transformer methods, including DiT [49], SiT [44], and the most recent yet contemporaneous work REPA [70].
Table 1: Sampling quality on various datasets. We report the performance of GMem on CIFAR-10 (left), ImageNet $64\times 64$ (right), and ImageNet $256\times 256$ (middle). GMem achieves comparable FID with fewer sampling steps across multiple datasets, further highlighting its advantages in accelerating training.
CIFAR-10
METHOD NI ($\downarrow$) FID ($\downarrow$)
Traditional generative models
Diffusion GAN [67] 51.2M 3.75
Diffusion StyleGAN [65] 50M 3.19
Diffusion models (UNets)
DDPM [24] 50M 3.17
Score SDE (deep) [61] 166M 2.20
EDM [28] 200M 2.01
SiD [72] 800M 1.92
Diffusion (Transformer)
SiT-XL/2 [44] 25.6M 6.68
+ REPA [70] 10M 4.52
+ GMem (ours) 10M 1.59
+ GMem (ours) 2.6M 4.08
+ GMem (ours) 22.5M 1.22
ImageNet 64$\times$64
METHOD NI ($\downarrow$) FID ($\downarrow$)
Traditional generative models
DMD2 [69] 98M 1.28
StyleGAN-XL [57] - 1.52
Diffusion models (UNets)
SiD [72] 409M 1.52
CTM [30] 61M 1.73
EMD [28] 200M 2.20
DPM-Solver [43] - 3.42
EDM (Heun) [28] 2500.5M 2.44
EDM2 (Heun) [29] 2147.5M 1.33
Diffusion models (Transformer)
SiT-XL [44]
+ GMem (ours) 22.5M 2.10
ImageNet 256$\times$256
METHOD NI ($\downarrow$) FID ($\downarrow$)
Diffusion models (UNets)
ADM [12] 512M 10.94
VDM+++++ + [31] 1.43B 3.36
MDTv2-XL/2 [18] 332.8M 9.60
Simple diffusion [25] 1B 3.71
Diffusion models (Transformer)
MaskDiT [71] 1B 5.69
Simple diffusion [25] 1B 3.75
SD-DiT [73] 512M 5.06
DiT-XL/2 [49] 1.8B 9.60
SiT-XL/2 [44] 1.8B 8.30
+ REPA [70] 1B 5.90
+ GMem (ours) 100M 4.98
+ GMem (ours) 512M 3.56
Metrics.
In line with prior research [44, 25], all generation quality metrics are reported as FID-50K (FID[22] scores. For CIFAR-10 dataset, we use the training set as the reference set, and for ImageNet $64\times 64$ and ImageNet $256\times 256$, we use the validation set as the reference set. Following Karras et al. [28], we also report the Number of trained Images (NI) as a measure of training efficiency. With similar FID scores, the model with less NI is considered more efficient111 $100$M NI corresponds to a total of $\sim 400$K training steps with the batch size of $256$, and baselines of different batch-sizes or training steps can be converted to NI in a similar way. . Moreover, following [42], we evaluate the efficiency of the sampling by comparing the Number of Function Evaluations (NFE) required to achieve a certain FID score. Under our setting, a lower NFE corresponds to reduced computational costs during sampling, thus making the model more efficient.
Implementation details.
We strictly adhere to the setup details outlined in [44] to ensure a fair comparison. For latent-space generation, we utilize ImageNet [10], where images are pre-processed to a resolution of $256\times 256$. We follow the ADM framework [12] for additional preprocessing strategies. Followed by Yu et al. [70], the preprocessed images then been fed into a Stable Diffusion VAE [55] and embedded into the latent space $\mathbf{z}\in\mathbb{R}^{32\times 32\times 4}$. For model configurations, we employ the B/2, L/2, and XL/2 architectures, as introduced by SiT [44], for ImageNet $256\times 256$.
We consistently use the batch size of $128$ for CIFAR-10 following [61] and $256$ for ImageNet following [29] during training. Additional experimental implementation details are provided in Appendix C.
Sampler configurations.
Following SiT [44], we use Euler Maruyama sampler (sde solver) and set the NFE as $50$ for CIFAR-10 and $250$ for ImageNet $64\times 64$ and ImageNet $256\times 256$ by default.
5.2 Main Results
Table 2: Ablation study and sensitive analysis. All models are based on SiT-L/2 and trained for 100M NI (400K iterations under the batch size of 256) without using classifier-free guidance on ImageNet $256\times 256$. $\downarrow$ indicates lower values are better.
[…]
0.2 640,000 5 epochs SDE 5.81
0.2 128,000 5 epochs SDE 9.72
0.2 640,000 1 epoch SDE 5.85
0.2 1,280,000 (Full) 0 epoch SDE 4.21
0.1 640,000 5 epochs SDE 5.93
0.3 640,000 5 epochs SDE 5.68
0.2 640,000 5 epochs ODE 6.56
GMem surpasses SoTA methods in both training efficiency and effectiveness.
  1. (a)
    Pixel-space generation. For pixel-space generation, as shown in Table 1, we benchmark GMem against prior methods using FID and NI. GMem outperforms all diffusion transformer baselines and achieves competitive performance with traditional generative models, obtaining FID =1.22 on CIFAR-10 and FID =2.10 on ImageNet $64\times 64$.
  2. (b)
    Latent-space generation. For latent-space generation, in Table 1, we compare the performance and the training efficiency of GMem with other baselines on ImageNet $256\times 256$. We observe that GMem achieves a FID of 4.98 with merely $100$M NI, outperforming the current SoTA methods, including SiT and REPA. Moreover, we demonstrate that this result can be further optimized to 3.56 when extends the NI to $512$M.
GMem is a 46.7×46.7$\times$46.7 × training accelerator.
We compare the advantages of GMem in terms of training efficiency. Table 1 and Table 4 also presents the FID scores achieved by GMem with limited NI.
For the pixel-space generation, on CIFAR-10, GMem achieves performance similar to REPA with only $2.6$M NI, offering more than 3.85×3.85$\times$3.85 × speedup over REPA and over 10×10$\times$10 × speedup compared to traditional SiT.
In latent-space generation, GMem attains FID of 5.75 with NI$=64$M, surpassing the most efficient diffusion transformer baseline, REPA, which requires more than NI$=1000$M. This indicates a training speedup of over 16×. Furthermore, GMem achieves an FID of 7.26 at NI$=38.4$M, providing a speedup exceeding 46.7× compared to SiT, which achieves the similar FID $=8.3$ with NI$=1800$M.
GMem is a 10×10$\times$10 × sampling accelerator.
Table 3: FID with respect to NFE on ImageNet 256$\times$256. In this table, we present the impact of variations in NFE on FID. All models are based on SiT-L/2, trained for 100M NI (400K iterations with a batch size of 256) without using classifier-free guidance. GMem achieves higher image quality within $\texttt{NFE}=25$, demonstrating over a 10×10$\times$10 × improvement in sampling efficiency.
[…]
SiT-L/2 25.6M 250 18.8
+ REPA 25.6M 250 08.4
+ GMem 25.6M 250 5.8
+ GMem 25.6M 50 7.5
+ GMem 25.6M 25 12.3
As a benefit of substantially lower FID score, GMem is capable of generating competitive samples with significantly fewer NFEs, resulting in over a 5×5$\times$5 × in sampling efficiency. Table 5.2 presents the FID scores of GMem at various NFEs.
Specifically, reducing the NFE from $250$ to $50$ leads to increase in FID of $1.7$, while still outperforming the previous SoTA, REPA, thereby achieving a sampling speedup of more than 5×5$\times$5 ×. Additionally, decreasing the NFE from $250$ to $25$ allows GMem to generate samples that remain competitive with those of SiT, attains an FID of 12.3, providing more acceleration of more than 10×10$\times$10 ×.
GMem is NOT a memory machine.
Storing information related to source data often raises privacy and copyright concerns, particularly in the context of image generation [58, 64, 5]. Although our approach incorporates a memory bank $\mathbf{M}$, we clarify that GMem does not function as a system solely for storing pixel-level detailed information from images. Instead, the memory bank stores high-level semantic representations, enabling $\mathbf{M}$ to serve as an external knowledge repository. This design facilitates faster training and sampling processes while mitigating privacy considerations.
Figure 4: Diverse image generation. The figure presents selected samples from ImageNet $256\times 256$ generated by the SiT-XL/2 + GMem model using checkpoints with $\texttt{NI}=512M$ and without classifier-free guidance. The top row shows the original image corresponding to the memory snippet, followed by four rows of images generated through different noises when sampling. For instance, in the sixth column, GMem adeptly captures the semantic essence of a castle nightscape from memory snippets, while presenting variations in elements such as appearance, lighting, perspective, and other nuanced details.
We demonstrate the semantic information within the memory bank $\mathbf{M}$ through experiments conducted on ImageNet $256\times 256$. Specifically, for one memory snippet $\mathbf{s}$ in $\mathbf{M}$, we apply different noise during the sampling process and compare with the generated samples. As shown in Figure 4, GMem generates images that are semantically aligned with the original input while differ in specific details. This outcome confirm that GMem does not merely memorize $\mathbf{M}$ or learn a direct mapping from $\mathbf{M}$ to $\mathcal{Z}$, but instead effectively leverages the high-level semantic knowledge stored in $\mathbf{M}$ during the image generation process.
Appendix D further demonstrates GMem’s flexibility in capturing the knowledge from $\mathbf{M}$, where we interpolate between representations from two different classes to generate a series of smoothly transitioning images.
5.3 Ablation Study
The effectiveness of GMem depends on several key factors, including dropout ratio, bank size, bank training settings, solver, and backbone architectures. In this subsection, we set NI to $100M$ and conducted a series of ablation studies on ImageNet $256\times 256$. We systematically examined these factors and concluded that, although each variable has a slight impact on the optimal performance, GMem still consistently generates high-quality images efficiently. The results are reported in Table 5.2 and Table 4.
Dropout ratio of memory snippets.
We first investigate the impact of different dropout ratios on model performance. As shown in Table 5.2, introducing a moderate level of dropout proves advantageous, with a 20% rate applied to the memory snippets yielding the best results. We argue that this dropout mechanism implicitly reduces the model’s reliance on memory snippets, thus improving generalizability and overall performance.
Configurations of building memory bank.
We examine the impact of various memory bank configurations on performance, we reported the results in Table 5.2. Our findings indicate that larger memory banks generally yield superior results, with bank size $=1{,}280{,}000$ leading to optimal result of FID $=4.21$. We propose that excessively small memory banks impede the retention of essential knowledge, thereby degrading FID performance, while overly large memory banks increase storage requirements and computational costs. Typically, allocating 50% of total images in the dataset to the memory bank—approximately 1.8 GB for ImageNet $256\times 256$—provides satisfactory outcomes. Additionally, we demonstrate that training the memory bank for five epochs is sufficient to achieve the optimal results.
Choice of SDE/ODE solver.
We observe that SDE solver consistently produces higher quality images than the ODE solver, achieving a FID score reduction of 0.25, as shown in  Table 5.2. Consequently, we employ SDE solvers in all main experiments.
Figure 5: Ablation study on the variation of FID-50K with respect to NI. We analyze the impact of NI during the training of GMem using different backbones on CIFAR-10 dataset. Specifically, while GMem-XL initially shows a higher FID than GMem-L at the early stages of training, after $1.75$M NI, GMem-XL consistently outperforms both GMem-L and GMem-B in terms of FID and maintains this lower value until convergence, where the final FID reaches $1.22$.
Neural network architecture.
Table 4: FID comparisons with various SiT backbones. In this table, we report the FID of GMem with different backbones on ImageNet $256\times 256$ at various NI values. For comparison, we also present the performance of the state-of-the-art baseline REPA at similar NI or comparable FID values. Note that $\downarrow$ indicates that lower values are preferred and all results reported are without classifier-free guidance.
[…]
SiT-B/2 130M 100M 33.0
+ REPA 130M 100M 24.4
+ GMem 130M 100M 15.4
SiT-L/2 458M 100M 18.8
+ REPA 458M 180M 8.4
+ GMem 458M 100M 5.8
SiT-XL/2 675M 100M 17.2
+ REPA 675M 100M 7.9
+ GMem 675M 100M 5.0
SiT-XL/2 675M 1800M 8.3
+ GMem 675M 38.4M 7.3
+ REPA 675M 1024M 5.9
+ GMem 675M 64M 5.8
+ GMem 675M 512M 3.6
We delve into the scalability of the neural network architecture of GMem by varying its size. Table 4 presents the FID-50K scores of GMem across different model sizes on ImageNet $256\times 256$. Our findings reveal that larger models not only converge more rapidly but also achieve lower FID scores, aligning with the observations of diffusion transformers by Yu et al. [70] and Ma et al. [44]. Notably, compared to smaller models, GMem-XL consistently reaches the same FID levels more quickly and ultimately converges to a lower FID of 3.56 on ImageNet $256\times 256$.
This scalability trend extends to pixel-space generation. For example, on CIFAR-10, as depicted in  Figure 5, GMem-XL achieves an FID of 1.22 more efficiently than smaller model variants. These results underscore the scalability of GMem, demonstrating its ability to maintain high generation quality while enhancing training and sampling efficiency across diverse architectures and datasets.
5.4 Additional Applications
In this section, we present experimental results demonstrating that the Memory Bank employed by GMem exhibits both cross-dataset transferability and adaptability to downstream text-to-image (T2I) tasks.
Transferable memory bank across datasets.
Appendix E demonstrates the transferability and generalization of the memory bank used to guide GMem across different datasets. Specifically, we train $\textsc{GMem}_{\textrm{IN64}}$ and $\textsc{GMem}_{\textrm{CIFAR}}$ models on ImageNet $64\times 64$ and CIFAR-10, respectively, resulting in memory banks of $\mathbf{M}_{\textrm{IN64}}$ and $\mathbf{M}_{\textrm{CIFAR}}$. We then directly apply $\mathbf{M}_{\textrm{IN64}}$ to guide the sampling process of $\textsc{GMem}_{\textrm{CIFAR}}$. Our results show that $\textsc{GMem}_{\textrm{CIFAR}}$ is still able to generate information consistent with the knowledge provided by $\mathbf{M}_{\textrm{IN64}}$.
Application to T2I generation.
We further demonstrate that GMem can be effectively applied to T2I generation tasks by encoding text prompts into memory snippets $\hat{\mathbf{s}}$, which are subsequently utilized by GMem to generate corresponding images.
This integration is achieved through a two-step process:
  1. (a)
    Pretraining the mapping function: We pretrain a mapping function $\boldsymbol{\psi}$ that transforms text prompt distributions $\mathbf{p}({\text{text}})$ into memory snippet distributions $\mathbf{p}(\mathbf{s})$ using a simple contrastive loss [54].
  2. (b)
    Image generation: We then employ the pretrained mapping function $\boldsymbol{\psi}$ in conjunction with the trained GMem model $\boldsymbol{\theta}$ to generate images.
Comprehensive descriptions of the T2I generation process, experimental setup, and the resulting generated images are provided in Appendix E.
6 Conclusion
In this paper, we introduce a new paradigm for diffusion models, GMem, which accelerates both training and sampling processes by incorporating a memory bank. Our method achieves SoTA performance on CIFAR-10 and ImageNet $64\times 64$, improving training efficiency by more than 10$\times$ and sampling efficiency by over 5$\times$ across datasets. Furthermore, we demonstrate that our memory bank can be transferred across different datasets and can be easily adapted to T2I tasks by generating a memory bank, further showcasing the flexibility and generalization capability of GMem.
Supplementary Material
We include additional analysis, experiments details, and results of downstream tasks in this supplementary material. In Section A, we provide a detailed analysis and experimental settings of the memory bank. In Section B, we present additional experiment, including the ablation of different encoders. In Section C, we offer further implementation details on the training settings of GMem on the CIFAR-10 dataset, ImageNet 64$\times$64 dataset, and ImageNet 256$\times$256 dataset. In Section D, we provide detailed analysis on the interpolation experiments. In Section E, we demonstrate that the bank learned from one dataset can be easily transferred to another dataset without any additional training. Besides, in Section F, we also show that GMem can be seamlessly applied to text-to-image generation tasks. Finally, in Section G, we supplement the derivation of the objective functions for flow-based and diffusion-based models.
Appendix A Memory Bank
In this section we provide a detailed analysis of the memory bank, including the theoratic analysis, building details of the memory bank and the ablative study of the memory bank size related to performance on downstream tasks.
A.1 Analysis
To ensure that for every input $\mathbf{x}$ sampled from the data distribution $D$, the normalized feature $\nicefrac{{\boldsymbol{f}(\mathbf{x})}}{{\|\boldsymbol{f}(\mathbf{x})\|}}$ corresponds exactly to a vector $\mathbf{s}$ present in the matrix $\mathbf{M}$, we require that:
$\underset{\mathbf{s}\in\mathbf{M}}{\min}\left\{\|\mathbf{s}-\frac{\boldsymbol{f}(\mathbf{x})}{\|\boldsymbol{f}(\mathbf{x})\|}\|\right\}=0,\,\forall\mathbf{x}\sim D\,.$ (11)
Since we enforce $\|\mathbf{s}\|=1$, the condition in (11) implies that:
$\underset{\mathbf{s}\in\mathbf{M}}{\max}\left\{\|\mathbf{s}\cdot\frac{\boldsymbol{f}(\mathbf{x})}{\|\boldsymbol{f}(\mathbf{x})\|}\|\right\}=1,\,\forall\mathbf{x}\sim D\,.$ (12)
Consequently, our objective is to satisfy:
${\max}\left\{\mathbf{M}\frac{\boldsymbol{f}(\mathbf{x})}{\|\boldsymbol{f}(\mathbf{x})\|}\right\}=1,\,\forall\mathbf{x}\sim D\,.$ (13)
This leads us to formulate the optimization problem as follows:
$\operatorname*{arg\,min\,}_{\mathbf{M}\in\mathbb{R}^{n\times m}}\left\{\left(\mathbb{E}_{\mathbf{x}\sim D}\left[\max\left\{\mathbf{M}\frac{\boldsymbol{f}(\mathbf{x})}{\|\boldsymbol{f}(\mathbf{x})\|}\right\}-1\right]\right)^{2}\right\}\,.$ (14)
A.2 Setting
As described in Table 6 and Table 7, across all experiments, we set bank unit $m$ to 768 to match the size of the output hidden dimension of the encoder. Additional, we set the memory bank size $n$ to 50k for CIFAR-10 and 1.2M for ImageNet 64$\times$64 and ImageNet 256$\times$256.
When constructing the memory bank, we use the Euclidean distance as the similarity measure between the features output by encoder and the memory bank and use MSE as loss. We choose the Adam optimizer with a learning rate of 1e-2, $\beta_{1}$ of 0.9, and $\beta_{2}$ of 0.999. The memory bank is trained for 80 epochs on CIFAR-10 with the batch size of 64.
A.3 Building
We first detailed the building process of the memory bank. Initially, the memory bank is populated with features and then fine-tuned using the outputs from the encoder. Specifically, during fine-tuning, we compute batched features from both the encoder and the memory bank. We then calculate the similarity between these batched features and the snippets within the memory bank, and compute the loss based on the most similar snippet. Finally, we update the memory bank features according to the computed similarities.
Table 5: Bank ablation study. We also conduct an ablation study with respect to the bank size on CIFAR-10. Note that $\downarrow$ indicates that lower values are preferred.
[…]
SiT-XL/2
+ (GMem) 50K 22.4M 1.22
+ (GMem) 25K 22.4M 1.50
+ (GMem) 5K 22.4M 3.90
We present the results of our ablation study on memory bank size in  Table 5. Overall, a larger memory bank leads to better performance, as indicated by lower FID scores. Specifically, when the bank size is 50,000—equivalent to the entire dataset—the performance is optimal. Reducing the bank size to 25k results in a slight performance decrease to FID $=1.50$, yet it still surpasses the baseline. However, further reducing the bank size to 5k significantly degrades performance, indicating the importance of maintaining a relatively large memory bank for optimal results.
Overall, GMem requires less than 150MB of additional storage (the size of a 50K memory bank) to achieve a 10$\times$ faster training speed and a 5$\times$ faster inference speed on downstream tasks.
Appendix B Supplementary Experiments
In this seciton, we include a supplementary experiments that apply GMem to further validate the effectiveness of GMem on various downstream tasks.
B.1 Ablation Study on the Encoder
We conduct an ablation study to investigate the impact of different encoders on the performance of GMem. Specifically, we choose the visual encoder from CLIP [54] as an alternative to show that GMem can be applied across encoders.
Experimental Setup
To avoid redundant computational overhead, instead of training GMem from scratch, we train only a multilayer perceptron (MLP) with dimensions $768\times 768$ to map the features output by the CLIP visual encoder to the memory bank. We utilize the Adam optimizer with a learning rate of $1\times 10^{-4}$ and train the MLP for one epoch on the ImageNet $256\times 256$ dataset.
Results
We present the results of the ablation study in Figure 6.
Appendix C Implementation Details
C.1 Diffusion transformer architecture
We closely follow the architecture used in REPA [70] and SiT [44]. Similar to a Vision Transformer [13], In this architecture, the input image is divided into patches, reshaped into a one-dimensional sequence of length $N$, and then processed by the model. Unlike the original SiT, REPA includes additional modulation layers called AdaIN-zero layers at each attention block. These layers scale and shift each hidden state based on the given timestep and additional conditions.
For latent space generation, similar to REPA, our architecture uses a downsampled latent image $z=E(x)$ as input, where $x$ is an RGB image and $E$ is the encoder of the Stable Diffusion Variational Autoencoder (VAE) [55]. For pixel space generation, we remove the encoder and directly use the RGB image as input. Specifically, we modify the original SiT by changing the number of channels from 4 to 3 and directly feed the transformed RGB image into the model. For ImageNet $64\times 64$, as an example of pixel space generation, we aslo adjust the patch size of SiT from 4 to 2 to maintain a consistent sequence length of $N=256$ patches.
Furthermore, unlike the original REPA, instead of using class label embeddings, we utilize the CLS token from the target representation output by the encoder as an additional condition for the diffusion transformer.
Figure 6: Ablation study on the encoder: Image generation using the CLIP visual encoder with GMem applied to SiT-XL/2, trained on the ImageNet $256\times 256$ dataset.
C.2 Hyperprarameters
Additional implementation details
We implement our model based on the original REPA implementation [70]. We use the AdamW optimizer [33] with a constant learning rate of $1\times 10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$, without weight decay. To accelerate training, we employ mixed-precision (fp16) computation along with gradient clipping.
For latent space generation, we pre-compute compressed latent vectors from raw images using the Stable Diffusion VAE [55] and utilize these latent vectors as input. In contrast, for pixel space generation, we directly use the raw pixel data as input. Although we experimented with data augmentations such as flipping, we found that they did not significantly improve performance. Therefore, we do not apply any data augmentation in our experiments.
For projecting memory snippets into the backbone hidden dimension, we utilize a three-layer MLP with SiLU activations for diffusion transformers, following Yu et al. [70]. We provide detailed hyperparameter configurations in Tables 6 and 7.
Table 6: Training settings of CIFAR-10. We provide the training settings for all models and training algorithms on the CIFAR-10 dataset.
Model Size
B L XL
Model details
Batch size 128 128 128
Training iterations 200K 200k 200k
Learning rate 1e-4 1e-4 1e-4
Optimizer Adam Adam Adam
Adam $\beta_{1}$ 0.9 0.9 0.9
Adam $\beta_{2}$ 0.999 0.999 0.999
Interpolants
$\alpha_{t}$ $1-t$ $1-t$ $1-t$
$\sigma_{t}$ $t$ $t$ $t$
$\omega_{t}$ $\sigma_{t}$ $\sigma_{t}$ $\sigma_{t}$
Training Objective v-prediction v-prediction v-prediction
Sampler Euler Euler Euler
Sampling steps 50 50 50
Classifier-free Guidance $\times$ $\times$ $\times$
Training details of backbone
Capacity(Mparams) 130 458 675
Input dim. 32$\times$32$\times$3 32$\times$32$\times$3 32$\times$32$\times$3
Num. layers 12 24 28
Hidden dim. 768 1,024 1,152
Num. heads 12 12 16
Training details of GMem
Bank size 50k 50k 50k
Bank similarity measure Euclidean Euclidean Euclidean
Bank objective MSE MSE MSE
Encoder $f(\mathbf{x})$ DINOv2-B DINOv2-B DINOv2-B
Table 7: Training settings. We present the training settings for all models and training algorithms on the ImageNet 64$\times$64 dataset (left) and the ImageNet 256$\times$256 dataset (right).
Model Size
B L XL
Model details
Batch size 128 128 128
Training iterations 400K 400k 400k
Learning rate 1e-4 1e-4 1e-4
Optimizer Adam Adam Adam
Adam $\beta_{1}$ 0.9 0.9 0.9
Adam $\beta_{2}$ 0.999 0.999 0.999
Interpolants
$\alpha_{t}$ $1-t$ $1-t$ $1-t$
$\sigma_{t}$ $t$ $t$ $t$
$\omega_{t}$ $\sigma_{t}$ $\sigma_{t}$ $\sigma_{t}$
Training Objective v-prediction v-prediction v-prediction
Sampler Euler Euler Euler
Sampling steps 50 50 50
Classifier-free Guidance $\times$ $\times$ $\times$
Training details of backbone
Capacity (Mparams) 130 458 675
Input dim. 64$\times$64$\times$3 64$\times$64$\times$3 64$\times$64$\times$3
Num. layers 12 24 28
Hidden dim. 768 1,024 1,152
Num. heads 12 12 16
Training details of GMem
Bank size 1.2M 1.2M 1.2M
Bank similarity measure Euclidean Euclidean Euclidean
Bank objective MSE MSE MSE
Encoder $f(\mathbf{x})$ DINOv2-B DINOv2-B DINOv2-B
Model Size
B L XL
Model details
Batch size 256 256 256
Training iterations 700K 700k 700k
Learning rate 1e-4 1e-4 1e-4
Optimizer Adam Adam Adam
Adam $\beta_{1}$ 0.9 0.9 0.9
Adam $\beta_{2}$ 0.999 0.999 0.999
Interpolants
$\alpha_{t}$ $1-t$ $1-t$ $1-t$
$\sigma_{t}$ $t$ $t$ $t$
$\omega_{t}$ $\sigma_{t}$ $\sigma_{t}$ $\sigma_{t}$
Training Objective v-prediction v-prediction v-prediction
Sampler Euler Euler Euler
Sampling steps 50 50 50
Classifier-free Guidance $\times$ $\times$ $\times$
Training details of backbone
Capacity(Mparams) 130 458 675
Input dim. 32$\times$32$\times$3 32$\times$32$\times$3 32$\times$32$\times$3
Num. layers 12 24 28
Hidden dim. 768 1,024 1,152
Num. heads 12 12 16
Training details of GMem
Bank size 50k 50k 50k
Bank similarity measure Euclidean Euclidean Euclidean
Bank objective MSE MSE MSE
Encoder $f(\mathbf{x})$ DINOv2-B DINOv2-B DINOv2-B
Encoder
We use Dinov2-B [48] as the encoder across all experiments, as it has been shown to significantly enhance the learning of better representations in diffusion models [70]. Dinov2-B offers superior performance, making it an ideal choice for facilitating the efficient training for constructing the memory bank.
Memory Bank
For the memory bank size, we set it to 50K for CIFAR-10 and 1.2M for ImageNet 64$\times$64 and ImageNet 256$\times$256. Note it is essential to maintain a relatively large memory bank for optimal performance, as demonstrated in Section A. Further results on the impact of memory bank size on performance are provided in Section A, where we detailed how varying the memory bank size affects the model’s efficacy.
Computing Resources
All models are primarily trained on NVIDIA H800 8-GPU setups, each equipped with 80GB of memory. The training speed for GMem-XL is approximately 1.15 seconds per 1KNI.
Appendix D Interpolation
D.1 GMem is Not a Memory Machine
In this section, we provide further evidence that GMem is not merely a memory machine. Specifically, we demonstrate that GMem can generate high-quality samples even for a memory snippet $\hat{\mathbf{s}}$ that never appeared in the training set.
D.2 Interpolation Between Memory Snippets
To investigate this ability, we conduct an interpolation experiment. We use the ImageNet $256\times 256$ dataset and employ the model checkpoint obtained after 180M (NI), as described in  Table 7.
We randomly select two memory snippets $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ from the memory bank $\mathbf{M}$. We then create nine interpolated snippets $\hat{\mathbf{s}}_{i}$ by linearly interpolating between $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ with interpolation coefficients $\alpha_{i}$ ranging from 0.1 to 0.9 in increments of 0.1. The interpolated snippets are defined as:
$\hat{\mathbf{s}}_{i}=(1-\alpha_{i})\mathbf{s}_{1}+\alpha_{i}\mathbf{s}_{2},\quad\alpha_{i}=0.1i,\quad i=1,2,\ldots,9.$
Each interpolated memory snippet $\hat{\mathbf{s}}_{i}$ is then fed into the transformer block to generate images.
D.3 Interpolation results
Figure 7: Interpolation between memory snippets. The first and last columns show the original memory snippets $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$, respectively. The remaining columns show the generated images from the interpolated memory snippets $\hat{\mathbf{s}}_{i}$.
Figure 8: A more elaborate interpolation experiment. The first and last columns show the original memory snippets $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$, respectively. The remaining columns show the generated images from the interpolated memory snippets $\hat{\mathbf{s}}_{i}$. Different row stands for different noise applied when generating the images.
The results of this interpolation experiment are presented in Figure 7 and  Figure 8. We observe that the generated images from the interpolated memory snippets $\hat{\mathbf{s}}_{i}$ are of high quality and exhibit smooth transitions between the two original memory snippets $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$.
In the first row of  Figure 7, we interpolate between an ape and a dog. The dog’s face gradually transforms into a smoother visage, adapting to resemble the ape. This demonstrates representation space learned by GMem is semantically smooth. Surprisingly, in the second row, interpolating between a green snake and a long-faced dog results in a green reptilian creature that resembles both the snake and the dog. This indicates that when the model encounters unseen memory snippets, it can utilize the smooth latent space to generate images similar to those it has previously encountered.
The third and last rows showcase even more imaginative interpolations. Interpolating between a monkey and barbed wire results in an image of a monkey in a cage, while a dog and a red hat can be interpolated into a dog with a black gentleman’s hat. These outcomes suggest that the similarities captured by the model are not limited to visual resemblance but also encompass more abstract semantic similarities in the latent space.
We believe that this semantic similarity arises because our memory bank introduces additional semantic information, enabling the model to better understand the content of images. Consequently, the model generates images that align more closely with human intuition, rather than merely memorizing the images corresponding to each snippet.
Appendix E Transferability of the Memory Bank
In this section, we demonstrate the transferability of the Memory Bank across different models. Specifically, we show that the Memory Bank can be transferred between GMem models trained on different datasets. While applying a Memory Bank extracted from low-resolution images to high-resolution models (e.g., Latent Diffusion Models) may result in decreased image sharpness due to information bottlenecks, it can still enhance the diversity of the generated image.
E.1 Experimental Setup
To investigate the transferability, we trained a Memory Bank $\mathbf{M}_{\mathrm{CIFAR}}$ on the CIFAR-10 dataset and directly transferred it to a model trained on ImageNet 256 $\times$ 256 ${\textsc{GMem}}_{\mathrm{IN256}}$ to guide image generation. We used the checkpoint from the ImageNet model at 180M (NI) for generation. The detailed experimental settings are provided in  Table 6 and  Table 7.
E.2 Results
Figure 9: Transferability of the Memory Bank. Each row corresponding to a specifical class in CIFAR-10. Specifically, the class is from top to bottom: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck.
Figure 9 presents the generation results of our method on the ImageNet $256\times 256$ dataset. The images demonstrate that the transferred Memory Bank can effectively guide the high-resolution model. Though sharpness is limited due to information bottlenecks in memory snippets, it can improving the diversity of the generated images.
Appendix F Text-to-Image Generation
We propose that GMem can be seamlessly applied to text-to-image generation tasks. This extension only requires training a mapping from textual CLS features to the Memory Bank, which can be efficiently implemented using a simple multilayer perceptron (MLP).
During inference, we extract CLS features from the Text Encoder using class names as input. These features are mapped to Memory Bank snippets through a trained MLP, and then fed into the SiT backbone to generate the corresponding images. This method enables straightforward and effective generation of images from textual descriptions.
F.1 Experimental Setup
In our text-to-image experiments, we utilize BERT-base [11] as the text encoder, using class names as textual supervision signals and the CLS token as the textual feature representation. We employ the model trained on CIFAR-10 with 22.5M (NI) as the generator.
For mapping text features to the Memory Bank, we use a simple MLP with dimensions 768$\times$768. We train this mapping using the Adam optimizer with a learning rate of $1\times 10^{-4}$, optimizing the CLIP Loss [54] over 60 epochs.
F.2 Experimental Results
Figure 10: Text-to-image generation results. Each row corresponds to a specific class in CIFAR-10. Specifically, the classes are from top to bottom: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck.
We present the generated images in Figure 10. The results demonstrate that our method can effectively generate images that correspond closely to the provided textual descriptions.
Appendix G Derivation
In this paper, we delve into two types of generative models that learn the target distribution by training variants of denoising autoencoders: denoising diffusion probabilistic models (DDPM) and stochastic interpolants.
G.1 Denoising Diffusion Probabilistic Models
Diffusion models [24] aim to model a target distribution $p(\mathbf{x})$ by learning a gradual denoising process that transitions from a Gaussian distribution $\mathcal{N}(0,\mathbf{I})$ to $p(\mathbf{x})$. The core idea is to learn the reverse process $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ of a predefined forward process $q(\mathbf{x}_{t}|\mathbf{x}_{0})$, which incrementally adds Gaussian noise to the data starting from $\mathbf{x}_{0}\sim p(\mathbf{x})$ over $T$ time steps.
The forward process $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ is defined as:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\,\mathbf{x}_{0},\beta_{t}^{2}\mathbf{I}\right),$
where $\beta_{t}\in(0,1)$ are small, predefined hyperparameters.
In the DDPM framework introduced by Ho et al. [24], the reverse process $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is parameterized as:
$p(\mathbf{x}_{t-1}$ $|\mathbf{x}_{t})=$
$\mathcal{N}\Bigg{(}$ $\mathbf{x}_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\varepsilon_{\theta}(\mathbf{x}_{t},t)\right),\Sigma_{\theta}(\mathbf{x}_{t},t)\Bigg{)}$
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, $\varepsilon_{\theta}(\mathbf{x}_{t},t)$ is a neural network parameterized by $\theta$ and $\Sigma_{\theta}(\mathbf{x}_{t},t)$ represents the learned variance.
The model is trained using a simple denoising autoencoder objective:
$L_{\text{simple}}=\mathbb{E}_{\mathbf{x}_{0},\varepsilon,t}\left[\left\|\varepsilon-\varepsilon_{\theta}(\mathbf{x}_{t},t)\right\|_{2}^{2}\right],$
where $\varepsilon$ is sampled from a standard normal distribution and $t$ is uniformly sampled from $\{1,\dots,T\}$.
For the variance $\Sigma_{\theta}(\mathbf{x}_{t},t)$, Ho et al. [24] initially set it to $\sigma_{t}^{2}I$ with $\beta_{t}=\sigma_{t}^{2}$. However, Nichol and Dhariwal [47] demonstrated that performance improves when $\Sigma_{\theta}(\mathbf{x}_{t},t)$ is learned jointly with $\varepsilon_{\theta}(\mathbf{x}_{t},t)$. They propose optimizing the variational lower bound (VLB) objective:
$L_{\text{vlb}}=\exp\left(v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t}\right),$
where $v$ is a per-dimension component from the model output and $\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$.
By choosing a sufficiently large $T$ and an appropriate schedule for $\beta_{t}$, the distribution $p(\mathbf{x}_{T})$ approaches an isotropic Gaussian. This allows for sample generation by starting from random noise and iteratively applying the learned reverse process $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ to obtain a data sample $\mathbf{x}_{0}$ citepho2020denoising.
G.2 Stochastic interpolating
In contrast to DDPM, flow-based models [15, 41] address continuous time-dependent processes involving data samples $\mathbf{x}^{*}\sim p(\mathbf{x})$ and Gaussian noise $\varepsilon\sim\mathcal{N}(0,\mathbf{I})$ over the interval $t\in[0,1]$. The process is formulated as:
$\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\varepsilon,\quad\text{with}\quad\alpha_{0}=\sigma_{1}=1,\quad\alpha_{1}=\sigma_{0}=0,$
where $\alpha_{t}$ decreases and $\sigma_{t}$ increases as functions of $t$. There exists a probability flow ordinary differential equation (PF-ODE) characterized by a velocity field $\dot{\mathbf{x}}_{t}=\mathbf{v}(\mathbf{x}_{t},t)$, ensuring that the distribution at time $t$ matches the marginal $p_{t}(\mathbf{x})$.
The velocity $\mathbf{v}(\mathbf{x},t)$ is expressed as a combination of two conditional expectations:
$\mathbf{v}(\mathbf{x},t)=\mathbb{E}[\dot{\mathbf{x}}_{t}\mid\mathbf{x}_{t}=\mathbf{x}]=\dot{\alpha}_{t}\,\mathbb{E}[\mathbf{x}^{*}\mid\mathbf{x}_{t}=\mathbf{x}]+\dot{\sigma}_{t}\,\mathbb{E}[\varepsilon\mid\mathbf{x}_{t}=\mathbf{x}],$
which can be approximated by a model $v_{\theta}(\mathbf{x}_{t},t)$ through minimizing the training objective:
$L_{\text{velocity}}(\theta)=\mathbb{E}_{\mathbf{x}^{*},\varepsilon,t}\left[\left\|v_{\theta}(\mathbf{x}_{t},t)-\dot{\alpha}_{t}\mathbf{x}^{*}-\dot{\sigma}_{t}\varepsilon\right\|^{2}\right].$
This approach aligns with the reverse stochastic differential equation (SDE):
$d\mathbf{x}_{t}=\mathbf{v}(\mathbf{x}_{t},t)\,dt-\frac{1}{2}w_{t}s(\mathbf{x}_{t},t)\,dt+\sqrt{w_{t}}\,d\bar{\mathbf{W}}_{t},$
where the score function $s(\mathbf{x}_{t},t)$ is similarly defined as:
$s(\mathbf{x}_{t},t)=-\frac{1}{\sigma_{t}}\,\mathbb{E}[\varepsilon\mid\mathbf{x}_{t}=\mathbf{x}].$
To approximate $s(\mathbf{x}_{t},t)$, one can use a model $s_{\theta}(\mathbf{x}_{t},t)$ with the training objective:
$L_{\text{score}}(\theta)=\mathbb{E}_{\mathbf{x}^{*},\varepsilon,t}\left[\left\|\sigma_{t}s_{\theta}(\mathbf{x}_{t},t)+\varepsilon\right\|^{2}\right].$
Since $s(\mathbf{x},t)$ can be directly computed from $\mathbf{v}(\mathbf{x},t)$ for $t>0$ using the relation:
$s(\mathbf{x},t)=\frac{1}{\sigma_{t}}\cdot\frac{\alpha_{t}\mathbf{v}(\mathbf{x},t)-\dot{\alpha}_{t}\mathbf{x}}{\dot{\alpha}_{t}\sigma_{t}-\alpha_{t}\dot{\sigma}_{t}},$
it is sufficient to estimate either the velocity $\mathbf{v}(\mathbf{x},t)$ or the score $s(\mathbf{x},t)$.
According to Albergo et al. [1], stochastic interpolants satisfy the following conditions when $\alpha_{t}$ and $\sigma_{t}$ are chosen such that: 1. $\alpha_{t}^{2}+\sigma_{t}^{2}>0$ for all $t\in[0,1]$, 2. Both $\alpha_{t}$ and $\sigma_{t}$ are differentiable over the interval $[0,1]$, 3. Boundary conditions are met: $\alpha_{1}=\sigma_{0}=0$ and $\alpha_{0}=\sigma_{1}=1$.
These conditions ensure an unbiased interpolation between $\mathbf{x}_{0}$ and $\varepsilon$. Consequently, simple interpolants can be utilized by defining $\alpha_{t}$ and $\sigma_{t}$ as straightforward functions during training and inference. Examples include linear interpolants with $\alpha_{t}=1-t$ and $\sigma_{t}=t$, or variance-preserving (VP) interpolants with $\alpha_{t}=\cos\left(\frac{\pi}{2}t\right)$ and $\sigma_{t}=\sin\left(\frac{\pi}{2}t\right)$.
An additional advantage of stochastic interpolants is that the diffusion coefficient $w_{t}$ remains independent when training either the score or velocity models. This independence allows $w_{t}$ to be explicitly chosen after training during the sampling phase using the reverse SDE.
It’s noteworthy that existing score-based diffusion models, including DDPM [24], can be interpreted within an SDE framework. Specifically, their forward diffusion processes can be viewed as predefined (discretized) forward SDEs that converge to an equilibrium distribution $\mathcal{N}(0,\mathbf{I})$ as $t\to\infty$. Training is conducted over $[0,T]$ with a sufficiently large $T$ (e.g., $T=1000$) to ensure that $p(\mathbf{x}_{T})$ approximates an isotropic Gaussian. Generation involves solving the corresponding reverse SDE, starting from random Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$. In this context, $\alpha_{t}$, $\sigma_{t}$, and the diffusion coefficient $w_{t}$ are implicitly defined by the forward diffusion process, potentially leading to a complex design space in score-based diffusion models [28].

### v2

@@
-Generative Modeling with Explicit Memory
+GMem: A Modular Approach for Ultra-Efficient Generative Models
 Abstract
-Corresponding author.* Equal contribution. Peng is the project leader.
 Recent studies indicate that the denoising process in deep generative diffusion models implicitly learns and memorizes semantic information from the data distribution.
 These findings suggest that capturing more complex data distributions requires larger neural networks, leading to a substantial increase in computational demands, which in turn become the primary bottleneck in both training and inference of diffusion models.
-To this end, we introduce Generative Modeling with Explicit Memory (GMem), leveraging an external memory bank in both training and sampling phases of diffusion models.
-This approach preserves semantic information from data distributions, reducing reliance on neural network capacity for learning and generalizing across diverse datasets.
-The results are significant: our GMem enhances both training, sampling efficiency, and generation quality.
-For instance, on ImageNet at $256\times 256$ resolution, GMem accelerates SiT training by over 46.7×46.7$\times$46.7 ×, achieving the performance of a SiT model trained for $7M$ steps in fewer than $150K$ steps.
-Compared to the most efficient existing method, REPA [70], GMem still offers a 16×16$\times$16 × speedup, attaining an FID score of 5.75 within $250K$ steps, whereas REPA requires over $4M$ steps.
-Additionally, our method achieves state-of-the-art generation quality, with an FID score of 3.56 without classifier-free guidance on ImageNet $256\times 256$.
-Our code is available at https://github.com/LINs-lab/GMem.
-Figure 1: GMem significantly enhances the training and sampling efficiency of diffusion models on ImageNet $256\times 256$.
-We propose an external memory bank that enhances the semantic information storage, improving both training stability and sampling efficiency in diffusion models.
-Sub-figure (a) illustrates the key concept of GMem, where $\boldsymbol{\epsilon}$ represents the initial noise and $\mathbf{x}_{0}$ the generated samples.
-In GMem, we disentangle the generalization and memorization capabilities originally intertwined within diffusion models, assigning memorization to an external memory bank $\mathbf{M}$.
-This decoupling reduces computational and capacity overhead for memorizing internal semantic information, thereby accelerating the overall process.
+To this end, we introduce GMem: A Modular Approach for Ultra-Efficient Generative Models.
+Our approach GMem decouples the memory capacity from model and implements it as a separate, immutable memory set that preserves the essential semantic information in the data.
+The results are significant: GMem enhances both training, sampling efficiency, and diversity generation.
+This design on one hand reduces the reliance on network for memorize complex data distribution and thus enhancing both training and sampling efficiency.
+On ImageNet at $256\times 256$ resolution, GMem achieves a 50×50$\times$50 × training speedup compared to SiT, reaching FID $=7.66$ in fewer than $28$ epochs ($\sim$4444 hours training time) [^1] All training time measurements are obtained on an 8×8$\times$8 ×H800 GPU cluster., while SiT requires $1400$ epochs.
+Without classifier-free guidance, GMem achieves state-of-the-art (SoTA) performance FID $=1.53$ in $160$ epochs with only $\sim 20$ hours of training, outperforming LightningDiT which requires $800$ epochs and $\sim$95959595 hours to attain FID $=2.17$.
+Additionally, the memory bank supports training-free adaptation to new images not present in the training set. Our code is available at https://github.com/LINs-lab/GMem.
+Machine Learning, ICML
+[…]
+[…]
+Figure 1: GMem Significantly enhances training and sampling efficiency of diffusion models on ImageNet $256\times 256$.
+We propose decoupling memorization capabilities from the model by implementing a separate, immutable memory bank that preserves essential data information.
+Sub-figure (a) highlights the core concept of GMem, where $\boldsymbol{\epsilon}$ denotes input noise and $\mathbf{x}_{0}$ represents generated samples.
+In GMem, we disentangle generalization and memorization capabilities, assigning memorization to an external memory bank $\mathbf{M}$.
+This decoupling reduces computational and capacity overhead, thus accelerating the process.
 Sub-figure (b) demonstrates the training efficiency of GMem on ImageNet $256\times 256$.
-At an FID of $5.8$, GMem achieves over 16×16$\times$16 × speedup relative to REPA [70].
-At an FID of $7.3$, GMem achieves over 46.7×46.7$\times$46.7 × speedup relative to SiT [44].
-Sub-figure (c) illustrates the sampling efficiency of GMem on ImageNet $256\times 256$.
-At the same target level of FID, GMem requires 5×5$\times$5 × less NFEs compared to REPA and 10×10$\times$10 × less NFEs compared to SiT.
+At an FID $=4.86$, GMem achieves over 25×25$\times$25 × speedup compared to REPA (Yu et al., 2024).
+At an FID $=7.66$, it achieves over 50×50$\times$50 × speedup relative to SiT (Ma et al., 2024).
+Sub-figure (c) illustrates sampling efficiency.
+At the same FID target, GMem requires 5×5$\times$5 × fewer NFEs compared to REPA and 10×10$\times$10 × fewer NFEs compared to SiT.
 1 Introduction
-Deep generative models based on denoising [68, 24, 59, 47, 9, 40, 8, 37]—which prioritize the generation of high-quality, realistic data—have achieved notable success within the deep learning community.
-These methods demonstrate exceptional performance in complex tasks such as zero-shot text-to-image and video generation [50, 56, 15, 51, 4].
-However, training and sampling diffusion models suffer from the issue of high computational burden [28, 29], due to the necessity of using neural networks with larger capacities for better empirical performance [28, 29].
-In the meanwhile, a line of research examines the diffusion models from the perspective of representation learning [36, 66, 8, 46], where diffusion models can capture semantic features within their hidden layers, with advanced models producing even stronger representations [66].
-Yang and Wang [68] illustrate that diffusion models strike a balance between learning meaningful features and regularizing the model capacity.
+Deep generative models based on denoising (Yang & Wang, 2023; Ho et al., 2020; Song et al., 2020a; Nichol & Dhariwal, 2021; Choi et al., 2021; Li et al., 2024; Chen et al., 2024c; Li et al., 2023c)—which prioritize the generation of high-quality, realistic data—have achieved notable success within the deep learning community.
+These methods demonstrate exceptional performance in complex tasks such as zero-shot text-to-image and video generation (Podell et al., 2023; Saharia et al., 2022; Esser et al., 2024; Polyak et al., 2024; Brooks et al., 2024).
+However, training and sampling diffusion models suffer from the issue of high computational burden (Karras et al., 2022, 2024), due to the necessity of using neural networks with larger capacities for better empirical performance (Karras et al., 2022, 2024).
+In the meanwhile, a line of research examines the diffusion models from the perspective of representation learning (Li et al., 2023a; Xiang et al., 2023; Chen et al., 2024c; Mukhopadhyay et al., 2021), where diffusion models can capture semantic features within their hidden layers, with advanced models producing even stronger representations (Xiang et al., 2023).
+Yang & Wang (2023) illustrate that diffusion models strike a balance between learning meaningful features and regularizing the model capacity.
 Kadkhodaie et al.
-[27] further interpret the generalization behavior in diffusion models, indicating that diffusion models aim to learn and encapsulate semantic information from the training data, subsequently generalizing to the true data distribution.
-In this paper, we extend these insights [68, 27] and conjecture that the functions of diffusion models can be implicitly divided into two components: (i) memorization of semantic information and (ii) generalization to true data.
-The intrinsic generation procedure of diffusion model then can be explained as transforming input noisy samples into the internal semantic distribution of a model, followed by “decoding” it into the data distribution.
-Specifically,
+(2023) further interpret the generalization behavior in diffusion models, indicating that diffusion models aim to learn and encapsulate semantic information from the training data, subsequently generalizing to the true data distribution.
+In this paper, we build upon the insights from prior works (Yang & Wang, 2023; Kadkhodaie et al., 2023) and propose a novel conjecture: the functionality of diffusion models can be implicitly decomposed into two distinct components: (i) memorization of semantic information and (ii) generalization to the true data distribution.
+We further posit that the intrinsic generation process of diffusion models can be interpreted as a transformation of noisy input samples into the model’s internal semantic representation, which is subsequently “decoded” into the target data distribution.
+To elaborate,
   1. (a)
-    in the training stage, diffusion models require significant neural network parameters and computational resources to extract and memorize semantic information from noisy samples;
+    During the training stage, diffusion models rely on substantial neural network capacity and computational resources to extract and memorize semantic information from noisy samples;
   2. (b)
-    in the sampling/inference stage, models often remap Gaussian noise to the internal semantic distribution before retrieving the data distribution, rather than directly transforming noise to data.
-Building upon the aforementioned analyses, the efficiency challenges of training and sampling diffusion models stem from the need to learn and memorize the internal semantic information, resulting in higher computational cost and greater demands on model capacity.
-As a remedy, we propose GMem, a new paradigm that uses an external memory bank to accompany the diffusion models to approximate the true data distribution, therefore avoiding memorizing the data distributions in the neural network.
-Extensive experiments demonstrate the effectiveness and efficiency of our approach, on both diffusion transformers [44] and convolutional neural networks [24, 61, 28].
-For example, integrating our GMem with SiT on ImageNet $256\times 256$ only requires $150K$ training iterations to achieve an FID of $7.3$, which is 46.7×46.7$\times$46.7 × faster compared to the baseline SiTs [44] and 16×16$\times$16 × than REPA [70].
-Under the same number of training iterations as $400K$, our model enhances the sampling efficiency by 10×10$\times$10 ×, where our model requires only $25$ sampling steps to achieve an $\texttt{FID}=12.3$ while the vanilla SiT needs $250$ sampling steps.
-Ultimately, our approach achieves SoTA performance with an FID of $3.56$ after $2M$ training steps, without classifier-free guidance.
-We outline the main contributions of this paper as follows:
+    During the sampling/inference stage, the models typically remap Gaussian noise to the internal semantic distribution before reconstructing the data distribution, rather than directly transforming noise into data.
+Building upon the aforementioned analyses, the efficiency challenges of training and sampling diffusion models stem from the need to learn and memorize internal semantic information.
+This process inherently incurs high computational costs and imposes significant demands on model capacity.
+As a remedy, we propose GMem, a new paradigm that decouples the memorization capability from the neural network by storing semantic information in an external memory bank.
+This separation eliminates the need for the neural network itself to memorize complex data distributions, enabling significant improvements: (1) accelerated training and sampling by avoiding the direct fitting of intricate semantic distributions; (2) flexible adjustment of memorization capacity in a training-free manner; (3) enhanced generalization, allowing the generation of images beyond the training set; and (4) reduced storage overhead by offloading memorization to the memory bank.
+Together, GMem provides a scalable and efficient framework for generative tasks.
+Extensive experiments across diverse diffusion backbones (Ma et al., 2024; Yao & Wang, 2025) and image tokenizers (Yao & Wang, 2025; Chen et al., 2024a) validate the effectiveness and efficiency of GMem.
+On ImageNet $256\times 256$, GMem achieves SoTA performance FID $=1.53$ in $160$ epochs with only $\sim$20202020 hours training time without classifier-free guidance.
+Under the same number of training epochs as $80$, GMem enhances the sampling efficiency by 10×10$\times$10 ×, where GMem requires only $25$ sampling steps to achieve an FID $=12.3$ while the vanilla SiT needs $250$ sampling steps, yielding 10×10$\times$10 × sampling time savings.
+We outline the main contributions of this paper below:
   1. (a)
-    We propose that diffusion models operate through two fundamental perspectives: memorization and generalization.
-Utilizing neural networks for memorization incurs significant computational and model capacity burdens.
+    We propose a novel conjecture that interprets the functionality of diffusion models into two distinct components: (i) memorization of semantic information and (ii) generalization to the true data distribution.
   2. (b)
-    We present GMem, a simple yet extremely effective approach to construct an external memory bank that stores semantic information inherent in data distributions, alleviating the memorization burden on neural networks within the diffusion models.
+    We introduce GMem, a simple yet highly effective approach that decouples the memorization capability of diffusion models into an external memory bank.
+This separation alleviates the memorization burden on the neural networks and enhances generalization flexibility.
   3. (c)
-    Our framework significantly enhances efficiency and quality of the SoTA diffusion models, for both training and inference (see Figure 1 and Figure 2).
-Figure 2: Selected samples on ImageNet $256\times 256$.
-This figure showcases a selection of images generated by GMem.
-We utilized the SiT-XL/2 backbone and trained GMem for $2M$ steps without classifier-free guidance.
+    GMem significantly enhances efficiency and quality of the SoTA diffusion models, for both training and inference (see Figure 1 and Figure 2).
+Figure 2: Selected samples on ImageNet $512\times 512$ and $256\times 256$.
+This figure presents images generated by GMem under two experimental settings: (1) For ImageNet $256\times 256$, GMem was trained for $160$ epochs and sampled via Euler method (NFE $=100$), achieving an FID $=1.53$ without classifier-free guidance.
+(2) For ImageNet $512\times 512$, training extended to $400$ epochs with identical sampling settings, yielding FID $=1.89$.
 2 Related work
 Generative models.
-Generative models—including Generative Adversarial Networks (GANs) [19, 57, 67], Variational Autoencoders (VAEs) [32, 21], flow-based methods, and diffusion-based methods [24, 12, 45]—aim to learn the data distribution $p(\mathbf{x})$ and generate data through sampling, demonstrating remarkable performance in producing realistic samples [38].
-Recently, diffusion-based methods utilize stochastic interpolation to model the forward process and subsequently convert the Gaussian distribution back into the original image distribution during the reverse process, yielding realistic samples.
-Such methods have achieved SoTA results in deep generative modeling and are the focus of this study [45, 60, 14].
-Training diffusion-based models is computationally intensive and unstable [70, 60].
-Additionally, the generative process necessitates multi-step sampling, incurring substantial computational costs [42] Accelerating the training and sampling processes of diffusion models has emerged as a significant challenge, prompting various efforts in previous studies to address this issue.
+Generative models–—including Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Sauer et al., 2022; Xiao et al., 2021), Variational Autoencoders (VAEs) (Kingma, 2013; He et al., 2022), flow-based methods, and diffusion-based methods (Ho et al., 2020; Dhariwal & Nichol, 2021; Mittal et al., 2023)—–aim to learn the data distribution $p(\mathbf{x})$ and generate data through sampling, achieving remarkable performance in producing realistic samples (Li et al., 2023d).
+Recently, diffusion-based methods employ stochastic interpolation to model a forward process and then reverse the Gaussian distribution back to the original image space, generating realistic samples.
+These methods achieve SoTA results in deep generative modeling and are the focus of this study (Mittal et al., 2023; Song & Ermon, 2020; Durkan & Song, 2021).
+Diffusion models face computational challenges due to high training cost and instability (Yu et al., 2024; Song & Ermon, 2020) and high sampling costs from multi-step generation (Lu & Song, 2024), driving extensive research to accelerate both processes.
+For example, REPA (Yu et al., 2024) leverages external visual representations to speed up training.
+LightningDiT (Yao & Wang, 2025) accelerates training by aligning the latent space of vision tokenizer (i.e.
+VA-VAE) with pretrained vision encoder.
+GMem instead introduces a novel decomposition of diffusion models into generalization and memorization components, enabling significantly faster training.
 Diffusion modeling and representation learning.
-To overcome the instability and computational inefficiency of diffusion models, recent studies  [70, 17, 45] have turned to representation learning to enhance diffusion models.
-On one hand, diffusion models are capable of learning high-quality representations [70].
+To overcome the instability and computational inefficiency of diffusion models, recent studies (Yu et al., 2024; Fuest et al., 2024; Mittal et al., 2023) start to leverage representation learning to enhance diffusion models.
+On the one hand, diffusion models are capable of learning high-quality representations (Yu et al., 2024).
 For instance, Tang et al.
-[63] demonstrated that feature maps extracted from diffusion networks can establish correspondences between real images, indicating a strong alignment between the learned representations with actual image.
-Furthermore, Yang and Wang [68] conducted a detailed analysis of the trade-off between the quality of learned representations and the penalization of the optimal parameter spectrum.
-Conversely, well-trained representation models can improve performance and expedite the training of diffusion models.
-[45] accomplish this by adjusting the weighting function in the denoising score matching objective to enhance representation learning.
-REPA [70] introduces an alignment loss for intermediate layer representations, significantly accelerating the training process by over $17.5$ times.
-In general, these methods enhance diffusion modeling by employing regularization techniques, which implicitly integrate semantic knowledge from pre-trained representation models.
-Self-supervised learning in diffusion models.
-Self-supervised learning, as a powerful approach within representation learning that leverages unlabeled data to learn useful representations, has garnered increasing attention in the diffusion modeling community.
-Many recent works [39, 26, 2] have adopted self-supervised learning techniques to enhance the representation learning capabilities of diffusion models, thereby improving their performance and efficiency.
-In particular, Li et al.
-[39] proposed a novel diffusion training paradigm that introduces an additional diffusion model to generate a 1D feature map as a condition to guide the training of the generative model.
-However, the inclusion of an additional diffusion model introduces extra sampling during inference, which increases computational overhead and can introduce additional errors.
-In this work, differing from Li et al.
-[39], we directly utilize representation from ‘memory bank’ as condition.
-This straightforward strategy eliminates the need for an additional diffusion model, thereby enhancing representation learning while reducing computational costs.
-It significantly improves both the training speed and the effectiveness of the model.
-In this work, differing from Li et al.
-[39], we directly utilize representation from ‘memory bank’ as condition.
-This straightforward strategy not only eliminates the unnecessary computational burden but also serves as an effective self-supervised learning method that enabling the model to learn better representations.
-As a result, our method significantly enhances both the training speed and effectiveness of diffusion model.
-Text to image generation.
-Text-to-image generation seeks to synthesize realistic images from free-form textual prompts.
-Despite advancements, current text-to-image models continue to experience challenges with text-image misalignment [53, 35].
-Recent works [6, 16] have attempted to address this issue by intervening in cross-modal attention activations that influence the internal states of Stable Diffusion models.
-However, this approach increases computational complexity and may impede the learning of optimal representations.
-Indeed, these models often require several days to weeks for training.
+(2023) demonstrate that feature maps extracted from diffusion networks can establish correspondences between real images, indicating a strong alignment between the learned representations with actual image.
+Furthermore, Yang & Wang (2023) conduct a detailed analysis of the trade-off between the quality of learned representations and the penalization of the optimal parameter spectrum.
+On the other hand, well-trained representation models can improve performance and expedite the training of diffusion models.
+Mittal et al.
+(2023) accomplish this by adjusting the weighting function in the denoising score matching objective to enhance representation learning.
+REPA (Yu et al., 2024) introduces an alignment loss for intermediate layer representations, significantly accelerating the training process by over $17.5$ times.
+External representation-augmented diffusion models via retrieval and generation.
+Retrieval-Augmented Generation (RAG) enhances generation quality by integrating external knowledge.
+IC-GAN (Casanova et al., 2021) augments image generation by conditioning on neighborhood instances retrieved from the training dataset.
+However, using only training images limits generalization.
+To address this issue, KNN-Diffusion (Sheynin et al., 2022) and RDM (Blattmann et al., 2022) employ large external memory sets, guiding generation via kNN retrieval during training and inference.
+Similarly, Chen et al.
+(2022) and Li et al.
+(2022) leverage a set of text-image pairs with cross-modality retrieval, improving generation performance on rare images.
+Despite their advantages, RAG methods encounter two key challenges: (1) substantial storage demands for large memory sets, and (2) increased computational costs during retrieval.
+We further contend that over-reliance on training sets restricts generalization capabilities (Blattmann et al., 2022).
+By employing a masking strategy (see our Section 5.4), we mitigate this dependency without incurring additional computational or storage overhead, thereby improving both generalization and training/inference efficiency.
+Representations for generation augmentation can also be obtained from representation generators.
+For instance, RCG (Li et al., 2023b) employs a representation generator to produce 1D “memory snippets” to guide diffusion models.
+Although RCG reduces the need to store large-scale memory sets, it encounters two primary challenges: (1) the requirement for additional training and sampling processes for the representation generator, increasing computational demands; (2) the necessity to retrain the representation generator to incorporate knowledge of new classes or artistic style transfers, thereby limiting its generalization capability.
+To overcome these limitations, Section 5.3 presents an efficient, training-free method for incorporating additional knowledge into the memory bank.
 3 Preliminary
-In this section, we provide a concise introduction to the training and sampling processes of flow-based and diffusion-based models. See Appendix G for more details.
-Both diffusion-based and flow-based models [44] derive their training procedures from a deterministic $T$-step noising process applied to the original data, formalized as:
+In this section, we provide a concise introduction to the training and sampling processes of flow-based and diffusion-based models. See Appendix F for more details.
+Both diffusion-based (Ho et al., 2020; Dhariwal & Nichol, 2021) and flow-based models (Ma et al., 2024) derive their training procedures from a deterministic $T$-step noising process applied to the original data (Ma et al., 2024), formalized as:
 $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})\,,$ (1)
 where $\mathbf{x}_{t}$ represents the noisy data at time $t$, $\mathbf{x}_{0}\sim p(\mathbf{x})$ is a real data sample from the true distribution, $\alpha_{t}$ and $\sigma_{t}$ are time-dependent decreasing and increasing functions respectively satisfying $\alpha_{t}^{2}+\sigma_{t}^{2}=1$.
-According to (1), each marginal probability density $p_{t}(\mathbf{x}_{t})$ corresponds to the distribution of a Probability Flow Ordinary Differential Equation [61] (PF ODE) with velocity field $\mathbf{v}(\mathbf{x},t)$, defined as:
+According to (1), each marginal probability density $p_{t}(\mathbf{x}_{t})$ corresponds to the distribution of a Probability Flow Ordinary Differential Equation (Song et al., 2020b) (PF ODE) with velocity field $\mathbf{v}(\mathbf{x},t)$, defined as:
 $\mathbf{v}(\mathbf{x},t)=\dot{\alpha}_{t}\,\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t}=\mathbf{x}]-\dot{\sigma}_{t}\,\mathbb{E}[\boldsymbol{\epsilon}\mid\mathbf{x}_{t}=\mathbf{x}]\,,$ (2)
@@
 By solving this ODE starting from $\mathbf{x}_{T}=\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, we obtain the probability density function $p_{0}(\mathbf{x}_{0})$, which can be used to estimate the ground-truth data distribution $p(\mathbf{x})$.
-Alternatively, the aforementioned noise-adding process can be formalized as a Stochastic Differential Equation [61, 59] (SDE):
+Alternatively, the aforementioned noise-adding process can be formalized as a Stochastic Differential Equation (Song et al., 2020b, a) (SDE):
 $\mathrm{d}\mathbf{x}_{t}=\boldsymbol{m}(\mathbf{x}_{t},t)\,\mathrm{d}t+g(t)\,\mathrm{d}\mathbf{W}_{t}\,,$ (3)
-where $\mathbf{W}_{t}$ is a Wiener process [23], $\boldsymbol{m}(\mathbf{x}_{t},t)$ is the drift coefficient defined as $\boldsymbol{m}(\mathbf{x}_{t},t)=-\frac{1}{2}\beta(t)\mathbf{x}_{t}$, and $g(t)$ is the diffusion coefficient, set as $g(t)=\sqrt{\beta(t)}$ with $\beta(t)$ being a time-dependent positive function controlling the noise schedule.
+where $\mathbf{W}_{t}$ is a Wiener process (Hitsuda, 1968), $\boldsymbol{m}(\mathbf{x}_{t},t)$ is the drift coefficient defined as $\boldsymbol{m}(\mathbf{x}_{t},t)=-\frac{1}{2}\beta(t)\mathbf{x}_{t}$, and $g(t)$ is the diffusion coefficient, set as $g(t)=\sqrt{\beta(t)}$ with $\beta(t)$ being a time-dependent positive function controlling the noise schedule.
 The corresponding reverse process is represented by the reverse-time SDE:
@@
 $s(\mathbf{x}_{t},t)=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})=-\frac{1}{\sigma_{t}}\mathbb{E}\left[\boldsymbol{\epsilon}\mid\mathbf{x}_{t}=\mathbf{x}\right]\,.$ (5)
-By solving the reverse-time SDE (4), starting from the initial state $\mathbf{x}_{T}=\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, we can obtain $p_{0}(\mathbf{x}_{0})$, thereby estimating the true data distribution $p(\mathbf{x})$.
+By solving the reverse-time SDE in (4), starting from the initial state $\mathbf{x}_{T}=\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, we can obtain $p_{0}(\mathbf{x}_{0})$, thereby estimating the true data distribution $p(\mathbf{x})$.
 4 Methodology
 Algorithm 1 Training GMem using memory bank $\mathbf{M}$
-procedure Train GMem($\mathbf{v}_{\boldsymbol{\theta}},\mathcal{D},\mathbf{M},T,\alpha_{t},\sigma_{t}$)
-     Initialize model parameters $\boldsymbol{\theta}$
-     for each training iteration do
-         Sample a batch of data $\mathbf{x}_{0}\sim\mathcal{D}$
-         Sample timesteps $t\sim\{0,\dots,T\}$ uniformly
-         Generate noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$
-         Compute noisy data $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\boldsymbol{\epsilon}$
-         Sample memory snippet $\mathbf{s}\sim\mathbf{M}$ using (9)
-         Predict velocity $\mathbf{v}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\mathbf{s})$
-         Compute loss using (8) and (10)
-         Backpropagate and update $\boldsymbol{\theta}$ using Optimizer
-     end for
-end procedure
-The motivation for our method arises from the observation that diffusion models inherently encode two distinct capabilities: generalization and memorization of semantic information within data [70, 68, 66, 8].
-We propose a paradigm in diffusion models that separates generalization and memorization: neural networks are employed for generalization, while an external memory bank is used to store semantic information of data.
-Specifically, we find that conventional generative models [24, 60] typically use a neural network to achieve both generalization and memorization of true data distributions.
-However, the UNet and Transformer-based neural networks [24, 49] used in modern diffusion models have limited parameters.
-This makes it challenging for these networks to memorize distribution information, leading to two main issues [27]: (a) limited parameters reduce the ability to memorize extensive data distributions, and (b) memorization via parameter optimization incurs high computational and time costs.
-To address these challenges and alleviate the memorization burden on neural networks in diffusion models, we propose employing an external memory bank to serve as the memorization function, therefore significantly improving training efficiency and the models’ ability to capture distributions.
-The framework is depicted in Figure 3.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+GMem is motivated by the observation that diffusion models inherently encode two distinct capabilities: generalization and memorization of semantic information within data (Yu et al., 2024; Yang & Wang, 2023; Xiang et al., 2023; Chen et al., 2024c).
+We introduce a novel paradigm that decouples these functions: neural networks handle generalization, while an external memory bank stores semantic information.
+Current generative models (Ho et al., 2020; Song & Ermon, 2020) typically rely on one neural network to simultaneously achieve both generalization and memorization of data distributions.
+However, the capacity-constrained modern diffusion models, such as UNet and Transformer-based networks (Ho et al., 2020; Peebles & Xie, 2023), face two critical limitations (Kadkhodaie et al., 2023): (1) insufficient model parameters to memorize complex data distributions, and (2) computationally expensive parameter optimization for memorization.
+To overcome these limitations, we propose offloading the memorization function to an external memory bank, thereby enhancing training efficiency and distribution capture capabilities.
+The proposed framework is illustrated in Figure 3.
 4.1 External Memory Bank
-Inspired by previous studies [27], which demonstrated that generalization in diffusion models arises from geometry-adaptive harmonic representations, we aim to design a memory bank that: (a) Provides sufficient semantic and representational information to assist the diffusion model in generating high-quality, realistic images that align with the true data distribution.
-(b) Avoids excessive detailed or textual information to prevent overfitting to the training data and ensure robust generalization.
-We define the memory bank for diffusion models as a matrix $\mathbf{M}\in\mathbb{R}^{n\times m}$, encompassing $n$ memory snippets.
-Each snippet $\mathbf{s}$ contains $m$ units and satisfies $\|\mathbf{s}\|=1$:
-$\mathbf{M}=\begin{bmatrix}\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{n}\end{bmatrix}^{\top}\,.$ (6)
-We employ a robust representation model $\boldsymbol{f}$ under the condition that for every input $\mathbf{x}$ sampled from data distribution $D$, the produced normalized feature $\nicefrac{{\boldsymbol{f}(\mathbf{x})}}{{\|\boldsymbol{f}(\mathbf{x})\|}}$ always corresponds to a vector $\mathbf{s}$ that exists in the set $\mathbf{M}$.
-In conclusion, our goal is to ensure that all semantic information contained in $D$ is captured by $\mathbf{M}$.
-To achieve this goal, we formalize the optimization objective for $\mathbf{M}$ as follows (see our analysis in Appendix A.1):
-$\operatorname*{arg\,min\,}_{\mathbf{M}\in\mathbb{R}^{n\times m}}\left\{\left(\mathbb{E}_{\mathbf{x}\sim D}\left[\max\left\{\mathbf{M}\frac{\boldsymbol{f}(\mathbf{x})}{\|\boldsymbol{f}(\mathbf{x})\|}\right\}-1\right]\right)^{2}\right\}\,,$ (7)
-where $\max(\cdot)$ denotes the maximum value of a vector.
-Note that $n$ and $m$ are predefined hyperparameters balancing the memorization capacity of $\mathbf{M}$ against storage constraints.
-Specifically, increasing $n$ and $m$ enhances the memorization capacity of $\mathbf{M}$ but also increases storage requirements.
-Detailed settings and analyses of $n$ and $m$ are provided in Appendix A.2.
-Usage of representation models for building the memory bank.
-Our memory bank framework is adaptable to various representation models $\boldsymbol{f}$, each with distinct capabilities in capturing image information.
-This variability impacts the performance of the trained generative models.
-Therefore, in this study, we utilize self-supervised representation models based on the following reasoning:
+Building on Kadkhodaie et al.
+(2023)’s insight that diffusion models achieve generalization through geometry-adaptive harmonic representations, we design a memory bank for diffusion models to: (a) supply essential semantic information for generating high-quality, realistic, and distribution-aligned images, and (b) exclude excessive details to prevent overfitting to the training data while maintaining robust generalization.
+We formalize the memory bank as a matrix $\mathbf{M}\in\mathbb{R}^{n\times m}$, consisting of $n$ unit-norm memory snippets:
+$\mathbf{M}=\begin{bmatrix}\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{n}\end{bmatrix}^{\top}\,,\quad|\mathbf{s}_{i}|=1\,.$ (6)
+We employ a representation model $\boldsymbol{f}$ such that for any input $\mathbf{x}\sim D$, the normalized feature $\nicefrac{{\boldsymbol{f}(\mathbf{x})}}{{\|\boldsymbol{f}(\mathbf{x})\|}}$ corresponds to a vector $\mathbf{s}\in\mathbf{M}$.
+Through optimization, we ensure that $\mathbf{M}$ fully captures the semantic information of $D$.
+Representation models for memory bank construction.
+GMem supports diverse representation models $\boldsymbol{f}$, each varying in their ability to capture image information, which also directly influences generative model performance.
+We employ self-supervised representation models for two key reasons:
   1. (a)
-    Previous research [70] demonstrates that self-supervised models can expedite the training of diffusion models, which can be seamlessly integrated with a memory bank.
+    Yu et al. (2024) demonstrate that self-supervised models can expedite the training of diffusion models. It motivates to use it in GMem.
   2. (b)
-    Prior work [3, 74, 62] points out that self-supervised models capture semantic information more effectively than supervised alternatives.
-We rigorously assess the effectiveness of different representation models within our framework to build the memory bank like CLIP visual encoder [54] and BERT [11] (see Appendix A.3 and Appendix F).
+    Self-supervised models capture semantic information more effectively than supervised alternatives (Bordes et al., 2022; Zimmermann et al., 2021; Sun et al., 2024).
+We examine multiple models, including CLIP visual encoder (Radford et al., 2021), to construct the memory bank (see Appendix E).
+Reducing network dependency on memory bank.
+Directly using memory snippets to guide network generation can lead to overfitting, as the network may rely too heavily on the strong guidance provided, impairing its ability to generalize to unseen samples (Blattmann et al., 2022).
+To address this issue, we propose a feature-level random masking strategy.
+Instead of utilizing the complete memory snippets during training, we randomly mask a subset of dimensions by setting them to zero.
+For instance, given a memory snippet with 512 dimensions, we can randomly select 256 dimensions to mask.
+We will detail the performance of various masking strategies and ratios in Section 5.4 and discuss how to select the most appropriate masking strategy.
 Figure 3: Data generation via GMem-enhanced diffusion models.
@@
 4.2 Training Objective
-Our proposed memory bank is designed to supply semantic information about the data distribution, supporting both the training and inference phases of diffusion models.
-Therefore, we incorporate semantic information into diffusion models during the training phase by slightly modifying the diffusion model loss function as follows:
+Memory bank provides semantic information about the data distribution, aiding both training and inference phases in diffusion models.
+To integrate this, we adapt the training loss of diffusion models as follows:
 […]
-where $\mathbf{x}_{0}\sim D,\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),\dot{\alpha}_{t}=\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t},\dot{\sigma}_{t}=\frac{\mathrm{d}\sigma_{t}}{\mathrm{d}t},\lambda\sim p(\lambda)$ and $\mathbf{v}_{\boldsymbol{\theta}}$ is the velocity estimated by SiT backbone.
-This approach trains diffusion models using a interpolant way over various noise levels [44], incorporating an additional input snippet $\mathbf{s}$, which was selected from memory bank $\mathbf{M}$ using cosine similarity:
-$\mathbf{s}=\operatorname*{arg\,max\,}_{\mathbf{s}_{i}\in\mathbf{M}}\left\{\frac{\boldsymbol{f}(\mathbf{x}_{0})}{\|\boldsymbol{f}(\mathbf{x}_{0})\|}\cdot\mathbf{s}_{i}\right\}\,.$ (9)
-To ensure a fair comparison with the baselines, we use the exact same parameter settings of $\alpha_{t}$ and $\sigma_{t}$ as in previous studies [44, 70].
-For a detailed description of the training and generation processes, please refer to the framework outlined in Algorithm 1.
-Multiple usage of representation models for enhancing training diffusion models.
-A very recent work REPA [70] introduces a regularization term that leverages representation models within the diffusion model’s loss function to enhance training efficiency.
-Therefore, we can integrate this regularization with our modified diffusion loss (8):
-$\mathcal{L}(\boldsymbol{\theta})\leftarrow\mathcal{L}(\boldsymbol{\theta})+\mathcal{L}_{\textsc{REPA}}(\boldsymbol{\theta})$ (10)
-$\text{with}\;\;\mathcal{L}_{\textsc{REPA}}(\boldsymbol{\theta}):=-\mathbb{E}_{\mathbf{x}_{0},t}$ $\left[\operatorname{sim}\left(\boldsymbol{f}(\mathbf{x}_{0}),\boldsymbol{\psi}(\mathbf{h})\right)\right]\,,$
-where $\boldsymbol{f}(\mathbf{x}_{0})$ denotes the pretrained representation, and $\boldsymbol{\psi}(\mathbf{h})$ represents the projection of the intermediate output of diffusion transformer block $\mathbf{h}$ using a trainable projection head $\boldsymbol{\psi}$, implemented as a multilayer perceptron (MLP).
+where $\mathbf{x}_{0}\sim D,\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),\dot{\alpha}_{t}=\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t},\dot{\sigma}_{t}=\frac{\mathrm{d}\sigma_{t}}{\mathrm{d}t}$ and $\mathbf{v}_{\boldsymbol{\theta}}$ is the velocity estimated by SiT backbone.
+For fair comparison, we retain the same $\alpha_{t}$ and $\sigma_{t}$ settings as in previous work (Ma et al., 2024; Yu et al., 2024).
 4.3 Sampling with Memory Bank
-The sampling stage of unconditional generation in diffusion models [24, 42] aims to transform Gaussian noise into the true data distribution within an end-to-end framework.
-This iterative process employs SDE or ODE solvers (refer to Section 3) to progressively convert the noise into the desired data distribution.
-Our framework GMem with the memory bank also follows this end-to-end generation process, namely, feeding the noise into the neural network, together with the corresponding memory snippet $\mathbf{s}$.
-Specifically, our GMem transforms the noise distribution into a corresponding bank snippet index distribution (a.k.a., noise-to-index transformation) and samples indices from this distribution to obtain memory snippets.
-We propose a two-step transformation methodology below—namely (1) normal distribution to uniform distribution and (2) uniform distribution to index distribution—that rigorously converts Gaussian-distributed noise into the index distribution.
-Such a strategy allows the index distribution exhibiting two key characteristics: firstly, similar noise inputs should correspond to similar indices, thereby maintaining coherence in the selection process; secondly, each generated index should possess an equal probability of being selected during the training phase, promoting unbiased learning.
-Step 1: normal distribution to uniform distribution.
-The initial step leverages the Probability Integral Transform [52] to convert Gaussian-distributed noise into a uniformly distributed variable.
-Specifically, we begin with a high-dimensional noise vector $\boldsymbol{\epsilon}\in\mathbb{R}^{d}$ that follows a Gaussian distribution.
-By computing and scaling the sum of the vector, we obtain a one-dimensional Gaussian variable $\overline{\epsilon}=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}\epsilon_{i}$, which follows the standard normal distribution $\mathcal{N}(0,1)$.
-The cumulative distribution function of this Gaussian variable, denoted by $\Phi(\overline{\epsilon})$, transforms the Gaussian noise into a uniform distribution over the interval $[0,1]$ and can be mathematically expressed as:
-[…]
-[…]
-Step 2: uniform distribution to index distribution.
-Given that the memory bank size may be much smaller than the total number of images in the dataset, a direct uniform distribution may not accurately reflect the true index distribution.
-To address this, we first record the index distribution from the training set, denoted as $\mathbf{P}_{\text{index}}=[p_{1},\ldots,p_{i},\ldots,p_{n}]$, where each $p_{i}$ represents the probability of selecting the $i$-th snippet $\mathbf{s}_{i}$ during training.
-Utilizing inverse transform sampling [52] , we map the uniform variable to the index distribution by defining the variable $k$ as the smallest index $i$ for which the cumulative probability $\sum_{j=1}^{i}p_{j}$ exceeds the uniform variable:
-$k=\min\left\{i\in\{1,\ldots,n\}\ \bigg{|}\ \Phi(\overline{\epsilon})\leq\sum_{j=1}^{i}p_{j}\right\}\,.$
-Upon completion of aforementioned noise-to-index transformation, the original noise vector $\boldsymbol{\epsilon}$ and the selected memory snippet $\mathbf{s}_{k}$ are combined and fed into the diffusion transformer backbone.
-The backbone is then interfaces with ODE or SDE solvers [24, 59] to generate images.
+The sampling stage of diffusion models (Ho et al., 2020; Lu & Song, 2024) transforms Gaussian noise into the true data distribution through an end-to-end framework, utilizing SDE or ODE solvers (see Section 3) to iteratively refine the noise.
+Similarly, GMem, equipped with a memory bank, follows this process by feeding noise and corresponding memory snippet $\mathbf{s}$ into the neural network.
+However, akin to Blattmann et al.
+(2022); Casanova et al.
+(2021), the memory bank requires explicitly storage and retrieval during inference.
+For large-scale datasets like ImageNet, this results in prohibitive storage demands and retrieval costs.
+Additionally, memory snippets often exhibit significant redundancy (e.g., snippets from the same class are highly correlated), leading to unnecessary storage when all features are retained directly.
+To reduce redundancy in storing memory snippets and minimize storage, we propose an efficient storage strategy.
+Decomposing the memory bank into two compact matrices.
+Let $\mathbf{M}\in\mathbb{R}^{N\times d}$ represent $N$ memory snippets $\mathbf{s}$, each of dimension $d$.
+We first compute the mean vector $\boldsymbol{\mu}\in\mathbb{R}^{d}$ and centralize $\mathbf{M}$ as $\mathbf{M}_{\text{centered}}=\mathbf{M}-\boldsymbol{\mu}$.
+Applying SVD (Eckart & Young, 1936) to it yields $\mathbf{M}_{\text{centered}}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{N\times r}$, $\mathbf{\Sigma}\in\mathbb{R}^{r\times r}$, and $\mathbf{V}\in\mathbb{R}^{d\times r}$, with $r<\min(N,d)$ as the target latent dimension.
+By distributing the singular values, we form a coefficient matrix $\mathbf{C}=\mathbf{U}\,\mathbf{\Sigma}^{1/2}\in\mathbb{R}^{N\times r}$ and a basis matrix $\mathbf{B}=\mathbf{V}\,\mathbf{\Sigma}^{1/2}\in\mathbb{R}^{d\times r}$.
+This results in a low-rank approximation:
+$\mathbf{M}\leftarrow\mathbf{C}\,\mathbf{B}^{\top}\,,$ (8)
+reducing storage complexity from $\mathcal{O}(Nd)$ to $\mathcal{O}(Nr+dr)$.
+The matrix $\mathbf{B}$ acts as a compact, fixed basis encoding global structure, whereas $\mathbf{C}$ flexibly stores snippet-specific coefficients.
+During inference, retrieving a snippet involves looking up its coefficients in $\mathbf{C}$ and transforming them via $\mathbf{B}$.
+Specifically, the $i$-th memory snippet $\mathbf{s}_{i}$ can be reconstructed as:
+$\mathbf{s}_{i}=\mathbf{c}_{i}\mathbf{B}^{\top}+\boldsymbol{\mu}\,,$ (9)
+where $\mathbf{c}_{i}\in\mathbb{R}^{1\times r}$ is the coefficient vector corresponding to the $i$-th snippet from coefficient matrix $\mathbf{C}$.
+4.4 Modular Memory Bank Enables Generalization
+The core insight of GMem lies in the decoupling of memorization and generalization in a generative model, enabling generalization beyond training data.
+This is achieved through two approaches: (I) incorporating external knowledge by introducing novel images absent from the original dataset, and (II) manipulating internal knowledge by combining existing memory snippets into new compositions.
+While our memory bank provides compact storage and flexibility for new snippets integration, we note that advancing the modularity of the memory bank remains future work.
+Aspect I: external knowledge manipulation.
+To incorporate external knowledge outside the training dataset, we project the feature vector $\boldsymbol{f}(\mathbf{x}_{\text{new}})\in\mathbb{R}^{d}$ onto the existing coefficient matrix $\mathbf{C}$.
+We calculate the coefficients $\mathbf{c}_{\text{new}}\in\mathbb{R}^{r}$ of the centered feature vector by projecting it onto the basis matrix $\mathbf{B}$:
+$\mathbf{c}_{\text{new}}=\nicefrac{{(\boldsymbol{f}(\mathbf{x}_{\text{new}})-\boldsymbol{\mu})\,\mathbf{B}}}{{\mathbf{S}}}\,,$ (10)
+where $\mathbf{S}$ (the diagonal of $\mathbf{\Sigma}$) stores singular values [^2] These singular values are several floating-point numbers with negligible storage cost..
+By appending $\mathbf{c}_{\text{new}}$ to $\mathbf{C}$ , we expanding the memory bank with negligible overhead.
+This process seamlessly integrates new snippets not present in the training dataset, allowing the model to utilize the network’s generalization capabilities for generating new samples without additional training.
+Aspect II: internal knowledge manipulation.
+We generate new memory snippets by interpolating between existing ones. Given two snippets indexed by $i$ and $j$, we construct a new coefficient vector:
+$\mathbf{c}_{\text{new}}=\alpha\,\mathbf{c}_{i}+(1-\alpha)\,\mathbf{c}_{j},\quad\alpha\in[0,1]\,.$ (11)
+Appending $\mathbf{c}_{\text{new}}$ to $\mathbf{C}$ yields a latent interpolation in $\mathbf{M}$ without modifying $\mathbf{B}$.
+This approach enables training-free style transfer and compositional generalization by exploring linear paths between coefficients of different memory snippets, effectively creating novel samples from the internal knowledge encoded in $\mathbf{C}$.
+In Section 5.3, we demonstrate the effectiveness of both exact projection and interpolation for expanding the memory bank, enhancing the diversity and flexibility of generated results.
 5 Experiments
 5.1 Experimental Setting
-We list the experimental settings below (see more details in Appendix B).
+We list the setup below (see more details in Appendix A).
 Datasets.
-For pixel space generation, we evaluate GMem on CIFAR-10 [34] for its diversity in classes and widespread use in benchmarking image generation models.
-We also evaluate the scalability of GMem on ImageNet $64\times 64$ [10] to assess the performance on larger resolution generation.
-We also conduct experiments on ImageNet $256\times 256$ [10] to evaluate the ability of GMem to model the latent space distribution, a key approach in current image generation research [28, 70, 44, 49].
-Backbones.
-Similar to prior image generation works [70], we utilizes SiT [44] as backbone.
-This architecture represent most recent advances in diffusion and flow-based transformers [70], and are widely used in image generation tasks [20, 7, 70].
+For pixel-level generation, we test GMem on CIFAR-10 (Krizhevsky et al., 2009) due to its diverse classes and its popularity in benchmarking image generation.
+We then evaluate GMem on ImageNet $256\times 256$ (Deng et al., 2009) to examine how it models latent space distributions, which is a key focus in recent image generation research (Karras et al., 2022; Yu et al., 2024; Ma et al., 2024; Peebles & Xie, 2023).
+Finally, we assess the scalability of GMem to larger resolutions by conducting experiments on ImageNet $512\times 512$ (Deng et al., 2009).
+Backbones and visual tokenizers.
+Following prior image generation approaches (Yu et al., 2024), we primarily use LightningDiT (Yao & Wang, 2025) and SiT (Ma et al., 2024) as the backbone.
+We also evaluate the effectiveness of GMem on different visual tokenizer such as SD-VAE (Rombach et al., 2022), DC-AE (Chen et al., 2024a), and VA-VAE (Yao & Wang, 2025) in  Table 5.
+These architectures represent the latest advancements in diffusion and flow-based transformers (Yu et al., 2024) and are widely adopted in image generation tasks (Han et al., 2023; Chen et al., 2024b; Yu et al., 2024).
 Baselines.
 For a fair comparison, we compare to the SoTA image generation methods on both training efficiency and performance.
-Specifically, for pixel-space image generation, we consider the following three categories of baselines: First, we compare GMem with traditional generative models, including Diffusion GAN [67], Diffusion StyleGAN [65], DMD2 [69].
-We also compare the SoTA diffusion models with UNets, including DDPM [24], Score SDE [61], EDM [28], DPM-Solver [43], ADM [12], EDMv2 [29], CTM [30], SiD [72], EMD [28].
-Finally, we also compare to the SoTA flow-based transformer methods, including DiT [49], SiT [44], and the most recent yet contemporaneous work REPA [70].
+Specifically, for pixel-space image generation, we consider the following three categories of baselines: First, we compare GMem with traditional generative models, including Diffusion GAN (Xiao et al., 2021), Diffusion StyleGAN (Wang et al., 2022), DMD2 (Yin et al., 2024).
+We also compare the SoTA diffusion models with UNets, including DDPM (Ho et al., 2020), Score SDE (Song et al., 2020b), EDM (Karras et al., 2022), DPM-Solver (Lu et al., 2022), ADM (Dhariwal & Nichol, 2021), EDMv2 (Karras et al., 2024), CTM (Kim et al., 2023), SiD (Zhou et al., 2024), EMD (Karras et al., 2022).
+Finally, we also compare to the SoTA flow-based transformer methods, including DiT (Peebles & Xie, 2023), SiT (Ma et al., 2024), and the most recent yet contemporaneous work REPA (Yu et al., 2024).
 Table 1: Sampling quality on various datasets.
-We report the performance of GMem on CIFAR-10 (left), ImageNet $64\times 64$ (right), and ImageNet $256\times 256$ (middle).
+We report the performance of GMem on CIFAR-10 (left), ImageNet $256\times 256$ (middle) and ImageNet $512\times 512$ (right).
 GMem achieves comparable FID with fewer sampling steps across multiple datasets, further highlighting its advantages in accelerating training.
+All the results reported are w/o classifier-free guidance unless otherwise specified.
 CIFAR-10
-METHOD NI ($\downarrow$) FID ($\downarrow$)
+METHOD Epoch ($\downarrow$) FID ($\downarrow$)
 Traditional generative models
-Diffusion GAN [67] 51.2M 3.75
-Diffusion StyleGAN [65] 50M 3.19
+BigGAN (Brock, 2018) - 14.7
+StyleGAN2 (Karras et al., 2020) 128 8.32
 Diffusion models (UNets)
-DDPM [24] 50M 3.17
-Score SDE (deep) [61] 166M 2.20
-EDM [28] 200M 2.01
-SiD [72] 800M 1.92
+DDPM (Ho et al., 2020) 100 3.17
+DDIM (Song et al., 2020a) - 4.04
+Score SDE (deep) (Song et al., 2020b) 333 2.20
+EDM (Karras et al., 2022) 400 2.01
+Diffusion Style-GAN (Wang et al., 2022) - 3.19
+Diffusion GAN  (Xiao et al., 2021) 1024 3.75
 Diffusion (Transformer)
-SiT-XL/2 [44] 25.6M 6.68
-+ REPA [70] 10M 4.52
-+ GMem (ours) 10M 1.59
-+ GMem (ours) 2.6M 4.08
-+ GMem (ours) 22.5M 1.22
-ImageNet 64$\times$64
-METHOD NI ($\downarrow$) FID ($\downarrow$)
+SiT-XL/2 (Ma et al., 2024) 512 6.68
+  + REPA (Yu et al., 2024) 200 4.52
+  + GMem (ours) 52 4.08
+  + GMem (ours) 200 1.59
+  + GMem (ours) 450 1.22
+ImageNet 256$\times$256
+METHOD Epoch ($\downarrow$) FID ($\downarrow$)
 Traditional generative models
-DMD2 [69] 98M 1.28
-StyleGAN-XL [57] - 1.52
+BigGAN (Brock, 2018) - 6.96
+VQ-GAN (Esser et al., 2021) - 15.78
 Diffusion models (UNets)
-SiD [72] 409M 1.52
-CTM [30] 61M 1.73
-EMD [28] 200M 2.20
-DPM-Solver [43] - 3.42
-EDM (Heun) [28] 2500.5M 2.44
-EDM2 (Heun) [29] 2147.5M 1.33
+ADM (Dhariwal & Nichol, 2021) 400 10.94
 Diffusion models (Transformer)
-SiT-XL [44]
-+ GMem (ours) 22.5M 2.10
-ImageNet 256$\times$256
-METHOD NI ($\downarrow$) FID ($\downarrow$)
+MaskGIT (Chang et al., 2022) 300 6.18
+MAGVIT-v2 (Yu et al., 2023) 270 3.65
+SD-DiT (Zhu et al., 2024) 480 7.21
+DiT-XL/2 (Peebles & Xie, 2023) 1400 9.62
+SiT-XL/2 (Ma et al., 2024) 1400 8.30
+  + REPA (Yu et al., 2024) 782 5.90
+FasterDiT (Yao et al., 2024) 400 7.91
+LightningDiT (Yao & Wang, 2025) 800 2.17
+  + GMem (ours) 160 1.53
+ImageNet 512$\times$512
+METHOD Epoch ($\downarrow$) FID ($\downarrow$)
+Traditional generative models
+StyleGAN-XL (Sauer et al., 2022) - 2.41
+BigGAN (Brock, 2018) 472 9.54
 Diffusion models (UNets)
-ADM [12] 512M 10.94
-VDM+++++ + [31] 1.43B 3.36
-MDTv2-XL/2 [18] 332.8M 9.60
-Simple diffusion [25] 1B 3.71
+ADM (Karras et al., 2024) - 23.2
+EDM2 (Karras et al., 2024) 734 1.91
 Diffusion models (Transformer)
-MaskDiT [71] 1B 5.69
-Simple diffusion [25] 1B 3.75
-SD-DiT [73] 512M 5.06
-DiT-XL/2 [49] 1.8B 9.60
-SiT-XL/2 [44] 1.8B 8.30
-+ REPA [70] 1B 5.90
-+ GMem (ours) 100M 4.98
-+ GMem (ours) 512M 3.56
+MaskGIT (Chang et al., 2022) 300 7.32
+MAGVIT-v2 (Yu et al., 2023) 270 3.07
+DiT-XL (Peebles & Xie, 2023) 600 12.03
+SiT-XL (Ma et al., 2024) (w/ cfg) 600 2.62
+  + REPA (Yu et al., 2024) (w/ cfg) 200 2.08
+LightningDiT-XL (Yao & Wang, 2025) - -
+  + GMem (ours) 200 1.93
+  + GMem (ours) 400 1.89
+  + GMem (ours) (w/ cfg) 400 1.71
 Metrics.
-In line with prior research [44, 25], all generation quality metrics are reported as FID-50K (FID[22] scores.
-For CIFAR-10 dataset, we use the training set as the reference set, and for ImageNet $64\times 64$ and ImageNet $256\times 256$, we use the validation set as the reference set.
-Following Karras et al.
-[28], we also report the Number of trained Images (NI) as a measure of training efficiency.
-With similar FID scores, the model with less NI is considered more efficient111 $100$M NI corresponds to a total of $\sim 400$K training steps with the batch size of $256$, and baselines of different batch-sizes or training steps can be converted to NI in a similar way.
-.
-Moreover, following [42], we evaluate the efficiency of the sampling by comparing the Number of Function Evaluations (NFE) required to achieve a certain FID score.
-Under our setting, a lower NFE corresponds to reduced computational costs during sampling, thus making the model more efficient.
+[…]
 Implementation details.
-We strictly adhere to the setup details outlined in [44] to ensure a fair comparison.
-For latent-space generation, we utilize ImageNet [10], where images are pre-processed to a resolution of $256\times 256$.
-We follow the ADM framework [12] for additional preprocessing strategies.
-Followed by Yu et al.
-[70], the preprocessed images then been fed into a Stable Diffusion VAE [55] and embedded into the latent space $\mathbf{z}\in\mathbb{R}^{32\times 32\times 4}$.
-For model configurations, we employ the B/2, L/2, and XL/2 architectures, as introduced by SiT [44], for ImageNet $256\times 256$.
-We consistently use the batch size of $128$ for CIFAR-10 following [61] and $256$ for ImageNet following [29] during training.
-Additional experimental implementation details are provided in Appendix C.
+Unless otherwise specified, we adhere to the setup details outlined in (Ma et al., 2024) to ensure a fair comparison.
+For latent-space generation, we utilize ImageNet $256\times 256$ and $512\times 512$.
+We follow the ADM framework (Dhariwal & Nichol, 2021) for additional preprocessing strategies.
+Following Ma et al.
+(2024), preprocessed images are fed into SD-VAE (Rombach et al., 2022) and embedded into the latent space $\mathbf{z}\in\mathbb{R}^{32\times 32\times 4}$.
+Additionally, we add representation alignment loss intrfoduced from REPA (Yu et al., 2024) to help the model learn more structured representations across all experiments.
+For model configurations, unless otherwise specified, we empolys SiT-XL for CIFAR-10 and LightningDiT-XL for ImageNet $256\times 256$ and $512\times 512$.
+We consistently use the batch size of $128$ for CIFAR-10 following (Song et al., 2020b) and $1024$ for ImageNet following (Karras et al., 2024) during training.
+Additional experimental implementation details are provided in Appendix B.
 Sampler configurations.
-Following SiT [44], we use Euler Maruyama sampler (sde solver) and set the NFE as $50$ for CIFAR-10 and $250$ for ImageNet $64\times 64$ and ImageNet $256\times 256$ by default.
-5.2 Main Results
-Table 2: Ablation study and sensitive analysis.
-All models are based on SiT-L/2 and trained for 100M NI (400K iterations under the batch size of 256) without using classifier-free guidance on ImageNet $256\times 256$.
-$\downarrow$ indicates lower values are better.
+Following SiT (Ma et al., 2024), we use SDE solver and set the NFE as $50$ for CIFAR-10 and $100$ for ImageNet $256\times 256$ and $512\times 512$ by default. More details are provided in Appendix B.
+Figure 4: Demonstration of novel and compositional image generation via memory bank manipulation.
+Selected samples from ImageNet $256\times 256$ generated by the GMem.
+In the “Novel image generation” part, we show the reference image used to build a new memory snippet (left), followed by the generated samples and 5 of the nearest training images, illustrating GMem’s adaptation to external knowledge.
+In the “Compositional image generation” examples, two reference images (left and right) form an interpolated image (center), demonstrating GMem can manipulate internal knowledge to create new concepts.
+Figure 5: Demonstration of diverse generation by GMem.
+Selected samples from ImageNet $256\times 256$ generated by the GMem.
+This figure demonstrates the diversity of images generated by GMem, which differ from the original training set in form, style, and color.
+This shows that GMem does not simply memorize images from the training set, but rather generates novel variations.
+5.2 Training and sampling efficiency of GMem
+Table 2: GMem offers 50$\times$ training speedup compared to SiT.
+This table compares training speedup provided by GMem to other baselines on ImageNet $256\times 256$.
+$\downarrow$ indicates lower is better, and all results are without classifier-free guidance.
 […]
-0.2 640,000 5 epochs SDE 5.81
-0.2 128,000 5 epochs SDE 9.72
-0.2 640,000 1 epoch SDE 5.85
-0.2 1,280,000 (Full) 0 epoch SDE 4.21
-0.1 640,000 5 epochs SDE 5.93
-0.3 640,000 5 epochs SDE 5.68
-0.2 640,000 5 epochs ODE 6.56
-GMem surpasses SoTA methods in both training efficiency and effectiveness.
-  1. (a)
-    Pixel-space generation.
-For pixel-space generation, as shown in Table 1, we benchmark GMem against prior methods using FID and NI.
-GMem outperforms all diffusion transformer baselines and achieves competitive performance with traditional generative models, obtaining FID =1.22 on CIFAR-10 and FID =2.10 on ImageNet $64\times 64$.
-  2. (b)
-    Latent-space generation.
-For latent-space generation, in Table 1, we compare the performance and the training efficiency of GMem with other baselines on ImageNet $256\times 256$.
-We observe that GMem achieves a FID of 4.98 with merely $100$M NI, outperforming the current SoTA methods, including SiT and REPA.
-Moreover, we demonstrate that this result can be further optimized to 3.56 when extends the NI to $512$M.
-GMem is a 46.7×46.7$\times$46.7 × training accelerator.
-We compare the advantages of GMem in terms of training efficiency. Table 1 and Table 4 also presents the FID scores achieved by GMem with limited NI.
-For the pixel-space generation, on CIFAR-10, GMem achieves performance similar to REPA with only $2.6$M NI, offering more than 3.85×3.85$\times$3.85 × speedup over REPA and over 10×10$\times$10 × speedup compared to traditional SiT.
-In latent-space generation, GMem attains FID of 5.75 with NI$=64$M, surpassing the most efficient diffusion transformer baseline, REPA, which requires more than NI$=1000$M.
-This indicates a training speedup of over 16×.
-Furthermore, GMem achieves an FID of 7.26 at NI$=38.4$M, providing a speedup exceeding 46.7× compared to SiT, which achieves the similar FID $=8.3$ with NI$=1800$M.
-GMem is a 10×10$\times$10 × sampling accelerator.
-Table 3: FID with respect to NFE on ImageNet 256$\times$256.
-In this table, we present the impact of variations in NFE on FID.
-All models are based on SiT-L/2, trained for 100M NI (400K iterations with a batch size of 256) without using classifier-free guidance.
-GMem achieves higher image quality within $\texttt{NFE}=25$, demonstrating over a 10×10$\times$10 × improvement in sampling efficiency.
+DiT-XL 675M 1400 9.62
+SiT-XL 675M 1400 8.61
+  + REPA 675M 800 5.90
+LightningDiT-XL 675M 800 2.17
+  + GMem 675M 28 7.66
+  + GMem 675M 32 4.86
+  + GMem 675M 160 1.53
+GMem is a $\mathbf{50\times}$ training accelerator.
+We evaluate training efficiency and performance across pixel-space and latent-space generation tasks.
+As shown in Table 1 and Table 5.2, GMem achieves competitive FID scores with significantly fewer training epochs.
+Pixel-space generation: On CIFAR-10 in Table 1, GMem matches REPA’s performance in just $450$ epochs, delivering a 3.85×3.85$\times$3.85 × speedup over REPA and over 10×10$\times$10 × compared to traditional SiT.
+Latent-space generation: On ImageNet in Table 1, GMem achieves FID $=1.53$ on ImageNet $256\times 256$ and $1.71$ on ImageNet $512\times 512$ in just $160$ and $400$ epochs, without using classifier-free guidance.
+As shown in Table 5.2, GMem achieves competitive FID scores with significantly fewer epochs than the most efficient baselines REPA and SiT: (1) With only $32$ epochs ($\mathbf{25\times}$ speedup), GMem achieves FID $=4.86$, outperforming REPA, which requires $800+$ epochs.
+(2) With only $28$ epochs ($\mathbf{50\times}$ speedup), GMem reaches FID $=7.66$, surpassing SiT’s FID $=8.61$ at $1{,}400$ epochs.
+GMem is a $\mathbf{10\times}$ sampling accelerator.
+Table 3: GMem accelerates sampling by 10$\times$.
+This table examines the impact of different ${NFE}$ choices on ${FID}$.
+For a fair comparison with REPA, we use SiT-L/2 as backbone and trained GMem for ${Epochs}=20$ (400K iterations with a batch size of 256) without using classifier-free guidance.
+$\downarrow$ means lower is better.
 […]
-SiT-L/2 25.6M 250 18.8
-+ REPA 25.6M 250 08.4
-+ GMem 25.6M 250 5.8
-+ GMem 25.6M 50 7.5
-+ GMem 25.6M 25 12.3
-As a benefit of substantially lower FID score, GMem is capable of generating competitive samples with significantly fewer NFEs, resulting in over a 5×5$\times$5 × in sampling efficiency.
-Table 5.2 presents the FID scores of GMem at various NFEs.
-Specifically, reducing the NFE from $250$ to $50$ leads to increase in FID of $1.7$, while still outperforming the previous SoTA, REPA, thereby achieving a sampling speedup of more than 5×5$\times$5 ×.
-Additionally, decreasing the NFE from $250$ to $25$ allows GMem to generate samples that remain competitive with those of SiT, attains an FID of 12.3, providing more acceleration of more than 10×10$\times$10 ×.
-GMem is NOT a memory machine.
-Storing information related to source data often raises privacy and copyright concerns, particularly in the context of image generation [58, 64, 5].
-Although our approach incorporates a memory bank $\mathbf{M}$, we clarify that GMem does not function as a system solely for storing pixel-level detailed information from images.
-Instead, the memory bank stores high-level semantic representations, enabling $\mathbf{M}$ to serve as an external knowledge repository.
-This design facilitates faster training and sampling processes while mitigating privacy considerations.
-Figure 4: Diverse image generation.
-The figure presents selected samples from ImageNet $256\times 256$ generated by the SiT-XL/2 + GMem model using checkpoints with $\texttt{NI}=512M$ and without classifier-free guidance.
-The top row shows the original image corresponding to the memory snippet, followed by four rows of images generated through different noises when sampling.
-For instance, in the sixth column, GMem adeptly captures the semantic essence of a castle nightscape from memory snippets, while presenting variations in elements such as appearance, lighting, perspective, and other nuanced details.
-We demonstrate the semantic information within the memory bank $\mathbf{M}$ through experiments conducted on ImageNet $256\times 256$.
-Specifically, for one memory snippet $\mathbf{s}$ in $\mathbf{M}$, we apply different noise during the sampling process and compare with the generated samples.
-As shown in Figure 4, GMem generates images that are semantically aligned with the original input while differ in specific details.
-This outcome confirm that GMem does not merely memorize $\mathbf{M}$ or learn a direct mapping from $\mathbf{M}$ to $\mathcal{Z}$, but instead effectively leverages the high-level semantic knowledge stored in $\mathbf{M}$ during the image generation process.
-Appendix D further demonstrates GMem’s flexibility in capturing the knowledge from $\mathbf{M}$, where we interpolate between representations from two different classes to generate a series of smoothly transitioning images.
+SiT-L/2 20 250 18.8
+  + REPA 20 250 08.4
+  + GMem 20 250 5.8
+  + GMem 20 50 7.5
+  + GMem 20 25 12.3
+By generating competitive samples with significantly fewer NFE, GMem improves sampling efficiency by over 5×5$\times$5 ×.
+Table 5.2 shows the FID scores of GMem across varying NFEs and baselines: (1) With NFE $=50$ ($\mathbf{5\times}$ speedup), GMem outperforms REPA with NFE $=250$.
+(2) With NFE only $25$ ($\mathbf{10\times}$ speedup), GMem still surpasses SiT with NFE $=250$.
+5.3 Manipulating Memory for Diverse Images
+Diverse image generation.
+We first demonstrate in Figure Figure 5 that even for memory snippets present in the training set, GMem can generate images that deviate significantly from the training data.
+This illustrates that GMem does not merely memorize images from the training set, but instead leverages the network’s learned generalization capabilities to produce novel samples.
+Specifically, when a memory snippet corresponding to the semantic concept of a "lighthouse" is provided, the generated image of the lighthouse exhibits clear differences in background, color, and other attributes compared to the most similar image in the training set.
+Similarly, when a memory snippet representing an "orange" is given, GMem generates an image of an orange placed on a plate, showing distinct differences in presentation from the close-up images of oranges in the training set.
+Novel image generation.
+The decoupled memory bank allows GMem to generate entirely new classes or samples without additional training.
+By integrating snippets via SVD-based methods (see Section 4.3), GMem can embed external knowledge into its memory bank and synthesize previously unseen categories.
+For instance, as shown in Figure 4, providing a “low-poly” style snippet lets GMem produce images exhibit “low-poly” characteristics but distinct from the original input—even if such a style was never in the training data.
+This training-free style transfer highlights how GMem can leverage memory snippets to create novel samples and underscores the flexibility of its decoupled memory bank in adapting to new styles or categories without retraining.
+Compositional image generation.
+Next, we will explore generating new concepts by manipulating existing snippets through latent space interpolation.
+Rather than introducing a new snippet, we linearly combine two snippets stored in $\mathbf{M}$ and the interpolated snippet produces images blending features from both “parent” snippets, creating hybrid or novel concepts.
+For example, interpolating between a “dog” and a “hat” snippet generates a new concept “dog wearing a hat”.
+This approach also allows introducing artistic styles to existing classes, e.g., we can generate a “swan swimming in a green river” by interpolating between a internal snippet “swan” and a external knowledge “green background” snippet.
+These results highlight GMem’s ability to composite internal and external knowledge for generating new concepts.
 5.4 Ablation Study
-The effectiveness of GMem depends on several key factors, including dropout ratio, bank size, bank training settings, solver, and backbone architectures.
-In this subsection, we set NI to $100M$ and conducted a series of ablation studies on ImageNet $256\times 256$.
-We systematically examined these factors and concluded that, although each variable has a slight impact on the optimal performance, GMem still consistently generates high-quality images efficiently.
-The results are reported in Table 5.2 and Table 4.
-Dropout ratio of memory snippets.
-We first investigate the impact of different dropout ratios on model performance.
-As shown in Table 5.2, introducing a moderate level of dropout proves advantageous, with a 20% rate applied to the memory snippets yielding the best results.
-We argue that this dropout mechanism implicitly reduces the model’s reliance on memory snippets, thus improving generalizability and overall performance.
-Configurations of building memory bank.
-We examine the impact of various memory bank configurations on performance, we reported the results in Table 5.2.
-Our findings indicate that larger memory banks generally yield superior results, with bank size $=1{,}280{,}000$ leading to optimal result of FID $=4.21$.
-We propose that excessively small memory banks impede the retention of essential knowledge, thereby degrading FID performance, while overly large memory banks increase storage requirements and computational costs.
-Typically, allocating 50% of total images in the dataset to the memory bank—approximately 1.8 GB for ImageNet $256\times 256$—provides satisfactory outcomes.
-Additionally, we demonstrate that training the memory bank for five epochs is sufficient to achieve the optimal results.
-Choice of SDE/ODE solver.
-We observe that SDE solver consistently produces higher quality images than the ODE solver, achieving a FID score reduction of 0.25, as shown in  Table 5.2.
-Consequently, we employ SDE solvers in all main experiments.
-Figure 5: Ablation study on the variation of FID-50K with respect to NI.
-We analyze the impact of NI during the training of GMem using different backbones on CIFAR-10 dataset.
-Specifically, while GMem-XL initially shows a higher FID than GMem-L at the early stages of training, after $1.75$M NI, GMem-XL consistently outperforms both GMem-L and GMem-B in terms of FID and maintains this lower value until convergence, where the final FID reaches $1.22$.
-Neural network architecture.
-Table 4: FID comparisons with various SiT backbones.
-In this table, we report the FID of GMem with different backbones on ImageNet $256\times 256$ at various NI values.
-For comparison, we also present the performance of the state-of-the-art baseline REPA at similar NI or comparable FID values.
-Note that $\downarrow$ indicates that lower values are preferred and all results reported are without classifier-free guidance.
-[…]
-SiT-B/2 130M 100M 33.0
-+ REPA 130M 100M 24.4
-+ GMem 130M 100M 15.4
-SiT-L/2 458M 100M 18.8
-+ REPA 458M 180M 8.4
-+ GMem 458M 100M 5.8
-SiT-XL/2 675M 100M 17.2
-+ REPA 675M 100M 7.9
-+ GMem 675M 100M 5.0
-SiT-XL/2 675M 1800M 8.3
-+ GMem 675M 38.4M 7.3
-+ REPA 675M 1024M 5.9
-+ GMem 675M 64M 5.8
-+ GMem 675M 512M 3.6
-We delve into the scalability of the neural network architecture of GMem by varying its size.
-Table 4 presents the FID-50K scores of GMem across different model sizes on ImageNet $256\times 256$.
-Our findings reveal that larger models not only converge more rapidly but also achieve lower FID scores, aligning with the observations of diffusion transformers by Yu et al.
-[70] and Ma et al.
-[44].
-Notably, compared to smaller models, GMem-XL consistently reaches the same FID levels more quickly and ultimately converges to a lower FID of 3.56 on ImageNet $256\times 256$.
-This scalability trend extends to pixel-space generation.
-For example, on CIFAR-10, as depicted in  Figure 5, GMem-XL achieves an FID of 1.22 more efficiently than smaller model variants.
-These results underscore the scalability of GMem, demonstrating its ability to maintain high generation quality while enhancing training and sampling efficiency across diverse architectures and datasets.
-5.4 Additional Applications
-In this section, we present experimental results demonstrating that the Memory Bank employed by GMem exhibits both cross-dataset transferability and adaptability to downstream text-to-image (T2I) tasks.
-Transferable memory bank across datasets.
-Appendix E demonstrates the transferability and generalization of the memory bank used to guide GMem across different datasets.
-Specifically, we train $\textsc{GMem}_{\textrm{IN64}}$ and $\textsc{GMem}_{\textrm{CIFAR}}$ models on ImageNet $64\times 64$ and CIFAR-10, respectively, resulting in memory banks of $\mathbf{M}_{\textrm{IN64}}$ and $\mathbf{M}_{\textrm{CIFAR}}$.
-We then directly apply $\mathbf{M}_{\textrm{IN64}}$ to guide the sampling process of $\textsc{GMem}_{\textrm{CIFAR}}$.
-Our results show that $\textsc{GMem}_{\textrm{CIFAR}}$ is still able to generate information consistent with the knowledge provided by $\mathbf{M}_{\textrm{IN64}}$.
-Application to T2I generation.
-We further demonstrate that GMem can be effectively applied to T2I generation tasks by encoding text prompts into memory snippets $\hat{\mathbf{s}}$, which are subsequently utilized by GMem to generate corresponding images.
-This integration is achieved through a two-step process:
-  1. (a)
-    Pretraining the mapping function: We pretrain a mapping function $\boldsymbol{\psi}$ that transforms text prompt distributions $\mathbf{p}({\text{text}})$ into memory snippet distributions $\mathbf{p}(\mathbf{s})$ using a simple contrastive loss [54].
-  2. (b)
-    Image generation: We then employ the pretrained mapping function $\boldsymbol{\psi}$ in conjunction with the trained GMem model $\boldsymbol{\theta}$ to generate images.
-Comprehensive descriptions of the T2I generation process, experimental setup, and the resulting generated images are provided in Appendix E.
+The performance of GMem depends on several factors: bank size, backbone architecture, solver, and masking strategies.
+We conduct ablation studies on ImageNet $256\times 256$ with epochs = $64$, and found that while each factor slightly impacts optimal performance, GMem consistently generates high-quality images efficiently.
+Optimal masking ratio for memory bank.
+We investigate the impact of varying masking ratio on model performance.
+As illustrated in Table A.2, a moderate mask ratio of $0.4$ balances performance and generalization, reducing reliance on memory snippets while avoiding performance degradation from excessive masking.
+This configuration achieves the best performance, yielding an FID $=5.70$.
+We also evaluate different masking strategies outlined in Appendix A.2. GMem adapts effectively to all three strategies, demonstrating robustness to noisy memory bank conditions.
+SVD-based decomposing strategy enhances generalization.
+We investigate the impact of memory bank compression methods, with results summarized in Table A.2.
+Applying the decomposing strategy from Section 4.3 while halving the memory bank size reduces FID by approximately $0.15$, likely due to increased sample diversity from slight noise.
+Additionally, we explore snippet-level compression by randomly selecting $640$ samples per class, resulting in a half-sized memory bank.
+This reduction degrades FID performance by only $0.02$, highlighting substantial redundancy in the memory bank and underscoring the necessity of compression.
+Additional findings.
+We show that SDE solver is always better than ODE solver in Appendix A.2 and Table A.2.
+We also find that larger backbones show better performance and GMem generalize well with various backbones and visual tokenizers in Appendix A.3 and Table 5.
 6 Conclusion
-In this paper, we introduce a new paradigm for diffusion models, GMem, which accelerates both training and sampling processes by incorporating a memory bank.
-Our method achieves SoTA performance on CIFAR-10 and ImageNet $64\times 64$, improving training efficiency by more than 10$\times$ and sampling efficiency by over 5$\times$ across datasets.
-Furthermore, we demonstrate that our memory bank can be transferred across different datasets and can be easily adapted to T2I tasks by generating a memory bank, further showcasing the flexibility and generalization capability of GMem.
-Supplementary Material
-We include additional analysis, experiments details, and results of downstream tasks in this supplementary material.
-In Section A, we provide a detailed analysis and experimental settings of the memory bank.
-In Section B, we present additional experiment, including the ablation of different encoders.
-In Section C, we offer further implementation details on the training settings of GMem on the CIFAR-10 dataset, ImageNet 64$\times$64 dataset, and ImageNet 256$\times$256 dataset.
-In Section D, we provide detailed analysis on the interpolation experiments.
-In Section E, we demonstrate that the bank learned from one dataset can be easily transferred to another dataset without any additional training.
-Besides, in Section F, we also show that GMem can be seamlessly applied to text-to-image generation tasks.
-Finally, in Section G, we supplement the derivation of the objective functions for flow-based and diffusion-based models.
-Appendix A Memory Bank
-In this section we provide a detailed analysis of the memory bank, including the theoratic analysis, building details of the memory bank and the ablative study of the memory bank size related to performance on downstream tasks.
-A.1 Analysis
-To ensure that for every input $\mathbf{x}$ sampled from the data distribution $D$, the normalized feature $\nicefrac{{\boldsymbol{f}(\mathbf{x})}}{{\|\boldsymbol{f}(\mathbf{x})\|}}$ corresponds exactly to a vector $\mathbf{s}$ present in the matrix $\mathbf{M}$, we require that:
-$\underset{\mathbf{s}\in\mathbf{M}}{\min}\left\{\|\mathbf{s}-\frac{\boldsymbol{f}(\mathbf{x})}{\|\boldsymbol{f}(\mathbf{x})\|}\|\right\}=0,\,\forall\mathbf{x}\sim D\,.$ (11)
-Since we enforce $\|\mathbf{s}\|=1$, the condition in (11) implies that:
-$\underset{\mathbf{s}\in\mathbf{M}}{\max}\left\{\|\mathbf{s}\cdot\frac{\boldsymbol{f}(\mathbf{x})}{\|\boldsymbol{f}(\mathbf{x})\|}\|\right\}=1,\,\forall\mathbf{x}\sim D\,.$ (12)
-Consequently, our objective is to satisfy:
-${\max}\left\{\mathbf{M}\frac{\boldsymbol{f}(\mathbf{x})}{\|\boldsymbol{f}(\mathbf{x})\|}\right\}=1,\,\forall\mathbf{x}\sim D\,.$ (13)
-This leads us to formulate the optimization problem as follows:
-$\operatorname*{arg\,min\,}_{\mathbf{M}\in\mathbb{R}^{n\times m}}\left\{\left(\mathbb{E}_{\mathbf{x}\sim D}\left[\max\left\{\mathbf{M}\frac{\boldsymbol{f}(\mathbf{x})}{\|\boldsymbol{f}(\mathbf{x})\|}\right\}-1\right]\right)^{2}\right\}\,.$ (14)
-A.2 Setting
-As described in Table 6 and Table 7, across all experiments, we set bank unit $m$ to 768 to match the size of the output hidden dimension of the encoder.
-Additional, we set the memory bank size $n$ to 50k for CIFAR-10 and 1.2M for ImageNet 64$\times$64 and ImageNet 256$\times$256.
-When constructing the memory bank, we use the Euclidean distance as the similarity measure between the features output by encoder and the memory bank and use MSE as loss.
-We choose the Adam optimizer with a learning rate of 1e-2, $\beta_{1}$ of 0.9, and $\beta_{2}$ of 0.999.
-The memory bank is trained for 80 epochs on CIFAR-10 with the batch size of 64.
-A.3 Building
-We first detailed the building process of the memory bank.
-Initially, the memory bank is populated with features and then fine-tuned using the outputs from the encoder.
-Specifically, during fine-tuning, we compute batched features from both the encoder and the memory bank.
-We then calculate the similarity between these batched features and the snippets within the memory bank, and compute the loss based on the most similar snippet.
-Finally, we update the memory bank features according to the computed similarities.
-Table 5: Bank ablation study. We also conduct an ablation study with respect to the bank size on CIFAR-10. Note that $\downarrow$ indicates that lower values are preferred.
-[…]
-SiT-XL/2
-+ (GMem) 50K 22.4M 1.22
-+ (GMem) 25K 22.4M 1.50
-+ (GMem) 5K 22.4M 3.90
-We present the results of our ablation study on memory bank size in  Table 5.
-Overall, a larger memory bank leads to better performance, as indicated by lower FID scores.
-Specifically, when the bank size is 50,000—equivalent to the entire dataset—the performance is optimal.
-Reducing the bank size to 25k results in a slight performance decrease to FID $=1.50$, yet it still surpasses the baseline.
-However, further reducing the bank size to 5k significantly degrades performance, indicating the importance of maintaining a relatively large memory bank for optimal results.
-Overall, GMem requires less than 150MB of additional storage (the size of a 50K memory bank) to achieve a 10$\times$ faster training speed and a 5$\times$ faster inference speed on downstream tasks.
-Appendix B Supplementary Experiments
+In this paper, we introduce a new paradigm for diffusion models, GMem, which accelerates both training and sampling processes by decoupling the memorization from the model backbone into a memory bank.
+GMem achieves SoTA performance on CIFAR-10 and ImageNet $256\times 256$ and $512\times 512$ and improving training efficiency by more than 50$\times$ and sampling efficiency by over 10$\times$ on ImageNet $256$.
+Additionally, the decoupled memory bank supports training-free adaptation to new images not present in the training set.
+Acknowledgements
+This work is supported in part by the Research Center for Industries of the Future (RCIF) at Westlake University, Westlake Education Foundation, and Westlake University Center for High-performance Computing.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix A Supplementary Experiments
 In this seciton, we include a supplementary experiments that apply GMem to further validate the effectiveness of GMem on various downstream tasks.
-B.1 Ablation Study on the Encoder
+A.1 Ablation Study on Vision Encoder
 We conduct an ablation study to investigate the impact of different encoders on the performance of GMem.
-Specifically, we choose the visual encoder from CLIP [54] as an alternative to show that GMem can be applied across encoders.
+Specifically, we choose the visual encoder from CLIP (Radford et al., 2021) as an alternative to show that GMem can be applied across encoders.
 Experimental Setup
@@
 We present the results of the ablation study in Figure 6.
-Appendix C Implementation Details
-C.1 Diffusion transformer architecture
-We closely follow the architecture used in REPA [70] and SiT [44].
-Similar to a Vision Transformer [13], In this architecture, the input image is divided into patches, reshaped into a one-dimensional sequence of length $N$, and then processed by the model.
+A.2 Ablation Study on Memory Bank Size, Masking Strategies and Solver
+Table 4: Ablation study and sensitive analysis.
+All models use FasterDiT-B, trained for ${Epochs}\!=\!64$ w/o using classifier-free guidance on ImageNet $256\times 256$.
+$\downarrow$ means lower is better.
+[…]
+1.2B Zero 0.4 SDE 5.70
+1.2B $\times$ Zero 0.4 SDE 5.85
+640K Zero 0.4 SDE 5.72
+1.2B Noise 0.4 SDE 6.79
+1.2B Random 0.4 SDE 6.62
+1.2B Zero 0 SDE 6.28
+1.2B Zero 0.3 SDE 5.75
+1.2B Zero 0.4 ODE 6.70
+Additional Masking Strategies.
+We also explored two other masking strategies: random mask and noise mask.
+Specifically, random mask replaces a randomly selected portion of each batch with Gaussian noise, while noise mask adds noise to the entire memory snippet.
+The results for these two masking strategies are presented in Table A.2.
+We found that zeroing out part of the snippet (the Zero mask strategy) consistently performed best across all experiments.
+Therefore, we adopted Zero mask for all major experiments.
+SDE solver is superior.
+SDE solvers consistently outperform ODE solvers, reducing FID by $1.0$ ( Table A.2). Thus, SDE solvers are used in all main experiments.
+A.3 Ablation Study on Model Backbones and Tokenizers
+Larger architectures excel.
+Table 5: GMem consistently generates high-quality samples across different backbones and image tokenizers.
+This table reports the FID of GMem with different backbones and visual tokenizers on ImageNet $256\times 256$.
+For a fair comparison, we train all models for 64 epochs.
+$\downarrow$ means lower is better and all results reported are without classifier-free guidance.
+[…]
+SiT-B SD-VAE 130M 64 22.25
+SiT-L SD-VAE 458M 64 6.49
+SiT-XL SD-VAE 675M 64 6.31
+LightningDiT-B VA-VAE 130M 64 5.70
+LightningDiT-B DC-AE 130M 64 5.97
+We examine the scalability of GMem by testing various model sizes and architectural configurations.
+Table 5 presents the FID-50K scores of GMem across different model sizes on ImageNet $256\times 256$: larger models not only converge faster but also achieve lower FID.
+This trend aligns with findings from Yu et al.
+(2024) and Ma et al.
+(2024) on diffusion transformers and extends to pixel-space generation.
+For example, on CIFAR-10, GMem-XL achieves an FID $=1.22$, outperforming smaller model variants efficiently, as depicted in  Figure 9.
+GMem generalize well with various backbones and visual tokenizers.
+We compare different backbone architectures and visual tokenizers in Figure 5.
+LightningDiT consistently outperforms SiT under identical configurations, corroborating findings from Yao & Wang (2025).
+Additionally, DC-AE and VA-VAE tokenizer yields better results than SD-VAE , likely due to their larger parameter capacity.
+Appendix B Implementation Details
+B.1 Diffusion transformer architecture
+We closely follow the architecture used in REPA (Yu et al., 2024) and SiT (Ma et al., 2024).
+Similar to a Vision Transformer (Dosovitskiy et al., 2021), In this architecture, the input image is divided into patches, reshaped into a one-dimensional sequence of length $N$, and then processed by the model.
 Unlike the original SiT, REPA includes additional modulation layers called AdaIN-zero layers at each attention block.
 These layers scale and shift each hidden state based on the given timestep and additional conditions.
-For latent space generation, similar to REPA, our architecture uses a downsampled latent image $z=E(x)$ as input, where $x$ is an RGB image and $E$ is the encoder of the Stable Diffusion Variational Autoencoder (VAE) [55].
+For latent space generation, similar to REPA, our architecture uses a downsampled latent image $z=E(x)$ as input, where $x$ is an RGB image and $E$ is the encoder of the Stable Diffusion Variational Autoencoder (VAE) (Rombach et al., 2022).
 For pixel space generation, we remove the encoder and directly use the RGB image as input.
@@
 Figure 6: Ablation study on the encoder: Image generation using the CLIP visual encoder with GMem applied to SiT-XL/2, trained on the ImageNet $256\times 256$ dataset.
-C.2 Hyperprarameters
+B.2 Hyperprarameters
 Additional implementation details
-We implement our model based on the original REPA implementation [70].
-We use the AdamW optimizer [33] with a constant learning rate of $1\times 10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$, without weight decay.
+We implement our model based on the original REPA implementation (Yu et al., 2024).
+We use the AdamW optimizer (Kingma & Ba, 2014) with a constant learning rate of $1\times 10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$, without weight decay.
 To accelerate training, we employ mixed-precision (fp16) computation along with gradient clipping.
-For latent space generation, we pre-compute compressed latent vectors from raw images using the Stable Diffusion VAE [55] and utilize these latent vectors as input.
+For latent space generation, we pre-compute compressed latent vectors from raw images using the Stable Diffusion VAE (Rombach et al., 2022) and utilize these latent vectors as input.
 In contrast, for pixel space generation, we directly use the raw pixel data as input.
@@
 For projecting memory snippets into the backbone hidden dimension, we utilize a three-layer MLP with SiLU activations for diffusion transformers, following Yu et al.
-[70].
+(2024).
 We provide detailed hyperparameter configurations in Tables 6 and 7.
@@
 Encoder $f(\mathbf{x})$ DINOv2-B DINOv2-B DINOv2-B
-Table 7: Training settings. We present the training settings for all models and training algorithms on the ImageNet 64$\times$64 dataset (left) and the ImageNet 256$\times$256 dataset (right).
+Table 7: Training settings. We present the training settings for all models and training algorithms on the ImageNet 256$\times$256 dataset (left) and the ImageNet 512$\times$512 dataset (right).
 Model Size
@@
 Model details
-Batch size 128 128 128
-Training iterations 400K 400k 400k
+Batch size 1024 1024 1024
+Training iterations 200K 200K 200K
 Learning rate 1e-4 1e-4 1e-4
@@
 Training Objective v-prediction v-prediction v-prediction
-Sampler Euler Euler Euler
-Sampling steps 50 50 50
-Classifier-free Guidance $\times$ $\times$ $\times$
+Sampler Heun Heun Heun
+Sampling steps 100 100 100
 Training details of backbone
 Capacity (Mparams) 130 458 675
-Input dim. 64$\times$64$\times$3 64$\times$64$\times$3 64$\times$64$\times$3
+Input dim. 32$\times$32$\times$3 32$\times$32$\times$3 64$\times$64$\times$3
 Num. layers 12 24 28
@@
 Bank size 1.2M 1.2M 1.2M
-Bank similarity measure Euclidean Euclidean Euclidean
-Bank objective MSE MSE MSE
 Encoder $f(\mathbf{x})$ DINOv2-B DINOv2-B DINOv2-B
@@
 Model details
-Batch size 256 256 256
-Training iterations 700K 700k 700k
+Batch size 1024 1024 1024
+Training iterations 500K 500k 500k
 Learning rate 1e-4 1e-4 1e-4
@@
 Training Objective v-prediction v-prediction v-prediction
-Sampler Euler Euler Euler
-Sampling steps 50 50 50
-Classifier-free Guidance $\times$ $\times$ $\times$
+Sampler Heun Heun Heun
+Sampling steps 100 100 100
 Training details of backbone
@@
 Training details of GMem
-Bank size 50k 50k 50k
-Bank similarity measure Euclidean Euclidean Euclidean
-Bank objective MSE MSE MSE
+Bank size 1.28M 1.28M 1.28M
 Encoder $f(\mathbf{x})$ DINOv2-B DINOv2-B DINOv2-B
 Encoder
-We use Dinov2-B [48] as the encoder across all experiments, as it has been shown to significantly enhance the learning of better representations in diffusion models [70].
+We use Dinov2-B (Oquab et al., 2023) as the encoder across all experiments, as it has been shown to significantly enhance the learning of better representations in diffusion models (Yu et al., 2024).
 Dinov2-B offers superior performance, making it an ideal choice for facilitating the efficient training for constructing the memory bank.
 Memory Bank
-For the memory bank size, we set it to 50K for CIFAR-10 and 1.2M for ImageNet 64$\times$64 and ImageNet 256$\times$256.
-Note it is essential to maintain a relatively large memory bank for optimal performance, as demonstrated in Section A.
-Further results on the impact of memory bank size on performance are provided in Section A, where we detailed how varying the memory bank size affects the model’s efficacy.
+For the memory bank size, we set it to 50K for CIFAR-10 and 1.2M for ImageNet 256$\times$256 and ImageNet 512$\times$512.
+Note it is essential to maintain a relatively large memory bank for optimal performance, as demonstrated in Section 5.4.
 Computing Resources
-All models are primarily trained on NVIDIA H800 8-GPU setups, each equipped with 80GB of memory. The training speed for GMem-XL is approximately 1.15 seconds per 1KNI.
-Appendix D Interpolation
-D.1 GMem is Not a Memory Machine
+All models are primarily trained on NVIDIA H800 8-GPU setups, each equipped with 80GB of memory. The training speed for GMem-XL is approximately 2.71 seconds per step.
+Appendix C Interpolation
+C.1 GMem is Not a Memory Machine
 In this section, we provide further evidence that GMem is not merely a memory machine.
 Specifically, we demonstrate that GMem can generate high-quality samples even for a memory snippet $\hat{\mathbf{s}}$ that never appeared in the training set.
-D.2 Interpolation Between Memory Snippets
-To investigate this ability, we conduct an interpolation experiment. We use the ImageNet $256\times 256$ dataset and employ the model checkpoint obtained after 180M (NI), as described in  Table 7.
+C.2 Interpolation Between Memory Snippets
+To investigate this ability, we conduct an interpolation experiment. We use the ImageNet $256\times 256$ dataset and employ the model checkpoint obtained after 140 epochs, as described in  Table 7.
 We randomly select two memory snippets $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ from the memory bank $\mathbf{M}$.
@@
 Each interpolated memory snippet $\hat{\mathbf{s}}_{i}$ is then fed into the transformer block to generate images.
-D.3 Interpolation results
+C.3 Interpolation results
 Figure 7: Interpolation between memory snippets.
@@
 Consequently, the model generates images that align more closely with human intuition, rather than merely memorizing the images corresponding to each snippet.
-Appendix E Transferability of the Memory Bank
+C.4 Additional Applications
+Figure 9: Ablation study on the variation of FID-50K with respect to number of training images (NI).
+We analyze the impact of NI during the training of GMem using different backbones on CIFAR-10 dataset.
+Specifically, while GMem-XL initially shows a higher FID than GMem-L at the early stages of training, after $1.75$M NI, GMem-XL consistently outperforms both GMem-L and GMem-B in terms of FID and maintains this lower value until convergence, where the final FID reaches $1.22$.
+In this section, we present experimental results demonstrating that the Memory Bank employed by GMem exhibits both cross-dataset transferability and adaptability to downstream text-to-image (T2I) tasks.
+Transferable memory bank across datasets.
+Appendix D demonstrates the transferability and generalization of the memory bank used to guide GMem across different datasets.
+Specifically, we train ${GMem}_{\textrm{IN64}}$ and ${GMem}_{\textrm{CIFAR}}$ models on ImageNet $64\times 64$ and CIFAR-10, respectively, resulting in memory banks of $\mathbf{M}_{\textrm{IN64}}$ and $\mathbf{M}_{\textrm{CIFAR}}$.
+We then directly apply $\mathbf{M}_{\textrm{IN64}}$ to guide the sampling process of ${GMem}_{\textrm{CIFAR}}$.
+Our results show that ${GMem}_{\textrm{CIFAR}}$ is still able to generate information consistent with the knowledge provided by $\mathbf{M}_{\textrm{IN64}}$.
+Application to T2I generation.
+We further demonstrate that GMem can be effectively applied to T2I generation tasks by encoding text prompts into memory snippets $\hat{\mathbf{s}}$, which are subsequently utilized by GMem to generate corresponding images.
+This integration is achieved through a two-step process:
+  1. (a)
+    Pretraining the mapping function: We pretrain a mapping function $\boldsymbol{\psi}$ that transforms text prompt distributions $\mathbf{p}({\text{text}})$ into memory snippet distributions $\mathbf{p}(\mathbf{s})$ using a simple contrastive loss (Radford et al., 2021).
+  2. (b)
+    Image generation: We then employ the pretrained mapping function $\boldsymbol{\psi}$ in conjunction with the trained GMem model $\boldsymbol{\theta}$ to generate images.
+Comprehensive descriptions of the T2I generation process, experimental setup, and the resulting generated images are provided in Appendix D.
+Appendix D Transferability of the Memory Bank
 In this section, we demonstrate the transferability of the Memory Bank across different models.
@@
 While applying a Memory Bank extracted from low-resolution images to high-resolution models (e.g., Latent Diffusion Models) may result in decreased image sharpness due to information bottlenecks, it can still enhance the diversity of the generated image.
-E.1 Experimental Setup
-To investigate the transferability, we trained a Memory Bank $\mathbf{M}_{\mathrm{CIFAR}}$ on the CIFAR-10 dataset and directly transferred it to a model trained on ImageNet 256 $\times$ 256 ${\textsc{GMem}}_{\mathrm{IN256}}$ to guide image generation.
-We used the checkpoint from the ImageNet model at 180M (NI) for generation.
+D.1 Experimental Setup
+To investigate the transferability, we trained a Memory Bank $\mathbf{M}_{\mathrm{CIFAR}}$ on the CIFAR-10 dataset and directly transferred it to a model trained on ImageNet 256 $\times$ 256 ${{GMem}}_{\mathrm{IN256}}$ to guide image generation.
+We used the checkpoint from the ImageNet model at 140 epochs for generation.
 The detailed experimental settings are provided in  Table 6 and  Table 7.
-E.2 Results
+D.2 Results
 Figure 10: Transferability of the Memory Bank.
@@
 Though sharpness is limited due to information bottlenecks in memory snippets, it can improving the diversity of the generated images.
-Appendix F Text-to-Image Generation
+Appendix E Text-to-Image Generation
 We propose that GMem can be seamlessly applied to text-to-image generation tasks.
@@
 This method enables straightforward and effective generation of images from textual descriptions.
-F.1 Experimental Setup
-In our text-to-image experiments, we utilize BERT-base [11] as the text encoder, using class names as textual supervision signals and the CLS token as the textual feature representation.
-We employ the model trained on CIFAR-10 with 22.5M (NI) as the generator.
+E.1 Experimental Setup
+In our text-to-image experiments, we utilize BERT-base (Devlin, 2018) as the text encoder, using class names as textual supervision signals and the CLS token as the textual feature representation.
+We employ the model trained on CIFAR-10 with 450 epochs as the generator.
 For mapping text features to the Memory Bank, we use a simple MLP with dimensions 768$\times$768.
-We train this mapping using the Adam optimizer with a learning rate of $1\times 10^{-4}$, optimizing the CLIP Loss [54] over 60 epochs.
-F.2 Experimental Results
+We train this mapping using the Adam optimizer with a learning rate of $1\times 10^{-4}$, optimizing the CLIP Loss (Radford et al., 2021) over 60 epochs.
+E.2 Experimental Results
 Figure 11: Text-to-image generation results.
@@
 We present the generated images in Figure 11. The results demonstrate that our method can effectively generate images that correspond closely to the provided textual descriptions.
-Appendix G Derivation
+Appendix F Derivation
 In this paper, we delve into two types of generative models that learn the target distribution by training variants of denoising autoencoders: denoising diffusion probabilistic models (DDPM) and stochastic interpolants.
-G.1 Denoising Diffusion Probabilistic Models
-Diffusion models [24] aim to model a target distribution $p(\mathbf{x})$ by learning a gradual denoising process that transitions from a Gaussian distribution $\mathcal{N}(0,\mathbf{I})$ to $p(\mathbf{x})$.
+F.1 Denoising Diffusion Probabilistic Models
+Diffusion models (Ho et al., 2020) aim to model a target distribution $p(\mathbf{x})$ by learning a gradual denoising process that transitions from a Gaussian distribution $\mathcal{N}(0,\mathbf{I})$ to $p(\mathbf{x})$.
 The core idea is to learn the reverse process $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ of a predefined forward process $q(\mathbf{x}_{t}|\mathbf{x}_{0})$, which incrementally adds Gaussian noise to the data starting from $\mathbf{x}_{0}\sim p(\mathbf{x})$ over $T$ time steps.
@@
 where $\beta_{t}\in(0,1)$ are small, predefined hyperparameters.
-In the DDPM framework introduced by Ho et al. [24], the reverse process $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is parameterized as:
+In the DDPM framework introduced by Ho et al. (2020), the reverse process $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is parameterized as:
 $p(\mathbf{x}_{t-1}$ $|\mathbf{x}_{t})=$
@@
 For the variance $\Sigma_{\theta}(\mathbf{x}_{t},t)$, Ho et al.
-[24] initially set it to $\sigma_{t}^{2}I$ with $\beta_{t}=\sigma_{t}^{2}$.
-However, Nichol and Dhariwal [47] demonstrated that performance improves when $\Sigma_{\theta}(\mathbf{x}_{t},t)$ is learned jointly with $\varepsilon_{\theta}(\mathbf{x}_{t},t)$.
+(2020) initially set it to $\sigma_{t}^{2}I$ with $\beta_{t}=\sigma_{t}^{2}$.
+However, Nichol & Dhariwal (2021) demonstrated that performance improves when $\Sigma_{\theta}(\mathbf{x}_{t},t)$ is learned jointly with $\varepsilon_{\theta}(\mathbf{x}_{t},t)$.
 They propose optimizing the variational lower bound (VLB) objective:
@@
 This allows for sample generation by starting from random noise and iteratively applying the learned reverse process $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ to obtain a data sample $\mathbf{x}_{0}$ citepho2020denoising.
-G.2 Stochastic interpolating
-In contrast to DDPM, flow-based models [15, 41] address continuous time-dependent processes involving data samples $\mathbf{x}^{*}\sim p(\mathbf{x})$ and Gaussian noise $\varepsilon\sim\mathcal{N}(0,\mathbf{I})$ over the interval $t\in[0,1]$.
+F.2 Stochastic interpolating
+In contrast to DDPM, flow-based models (Esser et al., 2024; Liu et al., 2023) address continuous time-dependent processes involving data samples $\mathbf{x}^{*}\sim p(\mathbf{x})$ and Gaussian noise $\varepsilon\sim\mathcal{N}(0,\mathbf{I})$ over the interval $t\in[0,1]$.
 The process is formulated as:
@@
 According to Albergo et al.
-[1], stochastic interpolants satisfy the following conditions when $\alpha_{t}$ and $\sigma_{t}$ are chosen such that: 1.
+(2023), stochastic interpolants satisfy the following conditions when $\alpha_{t}$ and $\sigma_{t}$ are chosen such that: 1.
 $\alpha_{t}^{2}+\sigma_{t}^{2}>0$ for all $t\in[0,1]$, 2.
@@
 This independence allows $w_{t}$ to be explicitly chosen after training during the sampling phase using the reverse SDE.
-It’s noteworthy that existing score-based diffusion models, including DDPM [24], can be interpreted within an SDE framework.
+It’s noteworthy that existing score-based diffusion models, including DDPM (Ho et al., 2020), can be interpreted within an SDE framework.
 Specifically, their forward diffusion processes can be viewed as predefined (discretized) forward SDEs that converge to an equilibrium distribution $\mathcal{N}(0,\mathbf{I})$ as $t\to\infty$.
@@
 Generation involves solving the corresponding reverse SDE, starting from random Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$.
-In this context, $\alpha_{t}$, $\sigma_{t}$, and the diffusion coefficient $w_{t}$ are implicitly defined by the forward diffusion process, potentially leading to a complex design space in score-based diffusion models [28].
+In this context, $\alpha_{t}$, $\sigma_{t}$, and the diffusion coefficient $w_{t}$ are implicitly defined by the forward diffusion process, potentially leading to a complex design space in score-based diffusion models (Karras et al., 2022).